### v1

# From the Choi Formalism in Infinite Dimensions to Unique Decompositions of Generators of Completely Positive Dynamical Semigroups

Abstract
Given any separable complex Hilbert space, any trace-class operator $B$ which does not have purely imaginary trace, and any generator $L$ of a norm-continuous one-parameter semigroup of completely positive maps we prove that there exists a unique bounded operator $K$ and a unique completely positive map $\Phi$ such that (i) $L=K(\cdot)+(\cdot)K^{*}+\Phi$, (ii) the superoperator $\Phi(B^{*}(\cdot)B)$ is trace class and has vanishing trace, and (iii) ${\rm tr}(B^{*}K)$ is a real number. Central to our proof is a modified version of the Choi formalism which relates completely positive maps to positive semi-definite operators. We characterize when this correspondence is injective and surjective, respectively, which in turn explains why the proof idea of our main result cannot extend to non-separable Hilbert spaces. In particular, we find examples of positive semi-definite operators which have empty pre-image under the Choi formalism as soon as the underlying Hilbert space is infinite-dimensional.
keywords:
Completely positive linear maps; Quantum channel; Completely positive dynamical semigroup; Choi matrix
MSC:
[2020] 37N20; 46N50; 47B10; 81P48
[…]
organization=Dahlem Center for Complex Quantum Systems, Freie Universität Berlin,addressline=Arnimallee 14, city=Berlin, postcode=14195, country=Germany
1 Introduction
Completely positive maps sit at the heart of quantum information theory and irreversible quantum dynamics, the latter of which captures fundamental physical processes such as decoherence or measurements. In particular, the evolution of many open quantum systems—that is, quantum systems which are not shielded from their environment—can be described by a norm-continuous one-parameter semigroup $(\Phi_{t})_{t\geq 0}$ of completely positive and trace-preserving maps on a complex Hilbert space $\mathcal{H}$; this is also known as quantum-dynamical semigroup (Breuer and Petruccione, 2002, Sec. 3.2). The semigroup property together with continuity in the parameter $t$ guarantees that the whole evolution is fully captured by the generator $L$ of the semigroup, that is, the (unique) bounded operator $L$ such that $\Phi_{t}=e^{tL}$ for all $t\geq 0$, refer to Engel and Nagel (2000) for more detail. What is more, there is even a standard form for generators of quantum-dynamical semigroups as first established in the seminal papers of Gorini et al. (1976) and Lindblad (1976): Every such $L$ can be written as111 Here, $[A,B]=AB-BA$ and $\{A,B\}=AB+BA$ are the usual commutator and anti-commutator, respectively. Moreover, $\Phi^{*}$ is the dual of $\Phi$ which we will recap properly at the end of Section 2. All that is important for now is that the specific choice of operator in the anti-commutator guarantees that the generated semigroup is trace-preserving. $-i[H,\cdot]+\Phi-\frac{1}{2}\{\Phi^{*}({\bf 1}),\cdot\}$ for some bounded, self-adjoint operator $H$ and some completely positive map $\Phi$; this is commonly known as GKSL-form. It should be noted that the difficult part here is to find conditions on the generator which guarantee complete positivity of the semigroup; in contrast, trace-preservation is simple as it merely amounts to the linear constraint ${\rm tr}(L(\rho))=0$ for all $\rho$. Thus, as a slight generalization one finds that $L$ generates a norm-continuous completely positive semigroup if and only if $L=K(\cdot)+(\cdot)K^{*}+\Phi$ for some $K\in\mathcal{B}(\mathcal{H})$ and some $\Phi$ completely positive (Christensen and Evans, 1979, Thm. 3.1).
From a physics perspective the term $-i[H,\cdot]$ in the generator represents the intrinsic evolution of the system (according to the Liouville-von Neumann equation) whereas $\Phi-\frac{1}{2}\{\Phi^{*}({\bf 1}),\cdot\}$ models the interaction of the system with its surroundings. Therefore, given some generator $L$, for applications and interpretation purposes it is desirable to know which part of the motion is due to the system itself and which part comes from the environment. The more precise question here would be whether there exist some “reasonable” domain and co-domain222 Note that such domain considerations are inevitable as for general $K,\Phi$ one can always shift terms the Kraus operators (more on those in Sec. 4) of $\Phi$ by a multiple of the identity which leads to a modification of $H$ while leaving the overall generator $L$ invariant (Davies, 1980, Eq. 1.4). such that the map $L\mapsto(H,\Phi)$ is well-defined. Indeed, this question is as old as the the GKSL-form itself: In their original work Gorini et al. (1976) have established that, for $\mathcal{H}$ finite-dimensional, such a unique decomposition is possible if both $H$ and $\Phi$ are traceless. While this condition has no meaningful counterpart in infinite dimensions, partial results in this direction have been achieved nonetheless: Uniqueness in infinite dimensions has been established for special classes of generators, cf. Alicki and Frigerio (1983), and for general $L$ under (rather restrictive) compactness conditions on the $K$ operator, cf. Davies (1980); Freiberger and Matthieu (1992). Moreover, it has been characterized when different $H$ and different Kraus operators (i.e. the “building blocks”) of $\Phi$ lead to the same generator $L$ (Parthasarathy, 1992, Thm. 30.16). Our main result will improve upon the result of Parthasarathy, the central point being that we turn equivalence classes of $(H,\Phi)$ into a unique decomposition by (i) fixing the Hamiltonian $H$ via a trace condition ${\rm tr}(B^{*}H)$ and, more importantly, by (ii) eliminating the ambiguity of the Kraus operators by replacing Parthasarathy’s trace condition (“${\rm tr}(B^{*}V_{j})=0$”) with a trace condition on the level of the completely positive map $\Phi$. Moreover, our result applies to all norm-continuous completely positive semigroups (as opposed to “just” quantum-dynamical semigroups) and the reference operators $B$ are as general as possible.
This paper’s main result builds upon a finding of ours for finite-dimensional spaces: Given any GKSL-generator $L$ and any $B\in\mathbb{C}^{n\times n}$ with ${\rm Re}({\rm tr}(B))\neq 0$ there exist unique $K\in\mathbb{C}^{n\times n}$ and unique $\Phi$ completely positive such that the map $X\mapsto\Phi(B^{*}XB)$ has zero trace, ${\rm Im}({\rm tr}(B^{*}K))=0$, and $L=K(\cdot)+(\cdot)K^{*}+\Phi$, cf. vom Ende (2023). Note that result by Gorini et al. is reproduced by setting $B={\bf 1}$. A key tool in proving the above result was the Choi matrix, that is, the matrix $\mathsf{C}(\Phi):=({\rm id}\otimes\Phi)(|\Gamma\rangle\langle\Gamma|)$ where $\Phi$ is an arbitrary linear map, $\Gamma:=\sum_{j=1}^{n}|j\rangle\otimes|j\rangle\in\mathbb{C}^{n}\otimes\mathbb{C}^{n}$ is the unnormalized entangled state, and $|j\rangle$ is the $j$-th standard basis vector. This formalism is used to relate completely positive maps to positive semi-definite operators on an enlarged Hilbert space (cf. Section 3 for more details). Now attempting to generalize the above result to infinite dimensions comes with a number of questions and challenges:
  • 1.
    Do the assumptions of the finite-dimensional result have meaningful counterparts in infinite dimensions? Of course, the reference matrix $B$ will become a trace-class operator, but it is not immediately clear when (resp. under which conditions) the trace of $\Phi(B^{*}(\cdot)B)$ even exists.
  • 2.
    What approaches to the Choi matrix in infinite dimensions exist, and which of them (if any) can help to prove our main result?
Moreover, even if these two questions were settled it is not yet clear whether (resp. under what conditions) the above result generalizes to infinite dimensions. We will address all of these problems in this work; this will not only lead to a generalization of vom Ende (2023) to arbitrary separable Hilbert spaces, but in the process we will obtain new results on the infinite-dimensional Choi formalism as well as on when linear maps between Schatten-class operators are themselves part of a Schatten class. This work is structured as follows. In Section 2 we recall some basic facts on Schatten classes as well as operators thereon. In Section 3 we review different approaches to the Choi formalism for general Hilbert spaces, and we characterize when the induced map that sends completely positive maps to positive semi-definite operators is injective and surjective (Proposition 1); the upshot there is that surjectivity cannot hold in infinite dimensions as certain reset maps $X\mapsto{\rm tr}(AX)B$ have empty preimage under the weighted Choi formalism. Then, in Section 4 we focus on Schatten-class operators which themselves act on Schatten classes; most importantly we find conditions on $\Phi$ and $B$ such that the map $X\mapsto\Phi(B^{*}XB)$ is trace class (Lemma 5). If $\Phi$ is completely positive, then the trace of $X\mapsto\Phi(B^{*}XB)$ can be computed explicitly via the Kraus operators of $\Phi$ (Lemma 6). Finally, Section 5 is dedicated to the generalization of vom Ende (2023)—that is, unique decompositions of generators of completely positive dynamical semigroups w.r.t. a reference operator $B$—to separable Hilbert spaces (Theorem 1). As a special case we obtain (a family of) unique decompositions for generators of quantum-dynamical semigroups (Corollary 1).
2 Preliminaries: Schatten Classes and Tensor Products
We begin by quickly recapping some (notation from) operator theory in general and Schatten classes in particular. Unless specified otherwise, $\mathcal{H},\mathcal{Z}$ will—here and henceforth—denote arbitrary complex Hilbert spaces. As is standard in mathematical physics we use the convention that the inner product is linear in the second variable. Moreover, we will frequently use bra-ket notation from quantum information theory: Given any $x\in\mathcal{Z},y\in\mathcal{H}$, $|x\rangle\langle y|:\mathcal{H}\to\mathcal{Z}$ is short for the linear operator $z\mapsto\langle y,z\rangle x$. This is also why, sometimes, we will write $|x\rangle$ instead of $x$. Be aware that as our Hilbert spaces may be non-separable we need to invoke nets as well as the concept of unordered summation in order to properly address questions of convergence333 Following Ringrose (1971)—or, alternatively, (Meise and Vogt, 1997, Ch. 12)—recall that a mapping $f:J\to\mathcal{X}$ from a non-empty set $J$ into a real or complex normed space $\mathcal{X}$ is called summable (to $x\in\mathcal{X}$) if the net $\{\sum_{j\in F}f(j)\}_{F\subseteq J\text{ finite}}$ norm-converges (to $x$). .
[…]
Lemma 1.
Given complex Hilbert spaces $\mathcal{H}_{1},\mathcal{H}_{2},\mathcal{H}_{3},\mathcal{H}_{4}$ and any $p,q,r>0$, the following statements hold.
  • (i)
    If $\frac{1}{r}=\frac{1}{p}+\frac{1}{q}$, then for all $X\in\mathcal{B}^{p}(\mathcal{H}_{2},\mathcal{H}_{3})$, $Y\in\mathcal{B}^{q}(\mathcal{H}_{1},\mathcal{H}_{2})$ one has $XY\in\mathcal{B}^{r}(\mathcal{H}_{1},\mathcal{H}_{3})$. Moreover, if $p,q,r\geq 1$, then $\|XY\|_{r}\leq\|X\|_{p}\|Y\|_{q}$.
  • (ii)
    If $\frac{1}{r}=\frac{1}{p}+\frac{1}{q}$, then for all $X\in\mathcal{B}^{r}(\mathcal{H}_{1},\mathcal{H}_{2})$ there exist $Y\in\mathcal{B}^{p}(\mathcal{H}_{2})$ and $Z\in\mathcal{B}^{q}(\mathcal{H}_{1},\mathcal{H}_{2})$ such that $X=YZ$.
  • (iii)
    If $q\geq p$, then for all $X\in\mathcal{B}^{p}(\mathcal{H}_{1},\mathcal{H}_{2})$, it holds that $\|X\|_{q}\leq\|X\|_{p}$. In particular, $\mathcal{B}^{p}(\mathcal{H}_{1},\mathcal{H}_{2})\subseteq\mathcal{B}^{q}(\mathcal{H}_{1},\mathcal{H}_{2})$ whenever $q\geq p$.
  • (iv)
    For all $X\in\mathcal{B}(\mathcal{H}_{3},\mathcal{H}_{4})$, $Z\in\mathcal{B}(\mathcal{H}_{1},\mathcal{H}_{2})$, $Y\in\mathcal{B}^{p}(\mathcal{H}_{2},\mathcal{H}_{3})$ one has $XYZ\in\mathcal{B}^{p}(\mathcal{H}_{1},\mathcal{H}_{4})$ with $\|XYZ\|_{p}\leq\|X\|\|Y\|_{p}\|Z\|$.
Proof.
We will only prove those statements which can not be found in the literature referred to above. (ii): By assumption $X$ admits a Schmidt decomposition $\sum_{j\in N}s_{j}(X)|f_{j}\rangle\langle g_{j}|$. Then $Y:=\sum_{j\in N}(s_{j}(X))^{r/p}|f_{j}\rangle\langle f_{j}|$ and $Z:=\sum_{j\in N}(s_{j}(X))^{r/q}|f_{j}\rangle\langle g_{j}|$ do the job. (iii): This result is stated in (Dunford and Schwartz, 1963, Ch. XI.9, Lemma 9) but without a proof, so we will fill this gap for the reader’s convenience. Let $X\in\mathcal{B}^{p}(\mathcal{H}_{1},\mathcal{H}_{2})$ and w.l.o.g. $X\neq 0$. We have to show $\|\,X\|X\|_{p}^{-1}\,\|_{q}\leq 1$. Defining $X^{\prime}:=X\|X\|_{p}^{-1}$ (i.e. $\|X^{\prime}\|_{p}=1$) one finds $s_{j}(X^{\prime})\leq 1$ for all $j\in N$; in particular this shows $(s_{j}(X^{\prime}))^{q/p}\leq s_{j}(X^{\prime})$ (as $\frac{q}{p}\geq 1$). This, as desired, implies $\|X^{\prime}\|_{q}\leq 1$ due to $\|X^{\prime}\|_{q}^{q}=\sum_{j\in N}(s_{j}(X^{\prime}))^{q}\leq\sum_{j\in N}(s_{j}(X^{\prime}))^{p}=1$. ∎
Important Schatten classes are the trace class $\mathcal{B}^{1}(\mathcal{H},\mathcal{Z})$—where, assuming $\mathcal{H}=\mathcal{Z}$, the trace ${\rm tr}(A):=\sum_{j\in J}\langle g_{j},Ag_{j}\rangle$ is well defined and independent of the chosen orthonormal basis $\{g_{j}\}_{j\in J}$ of $\mathcal{H}$—as well as the Hilbert-Schmidt class $\mathcal{B}^{2}(\mathcal{H},\mathcal{Z})$ which is itself a Hilbert space with respect to the inner product $\langle X,Y\rangle_{\sf HS}:={\rm tr}(X^{*}Y)$. A well-known, yet important fact is that an orthonormal basis of $\mathcal{B}^{2}(\mathcal{H},\mathcal{Z})$ is given by $\{|f_{k}\rangle\langle g_{j}|\}_{j\in J,k\in K}$ where $\{f_{k}\}_{k\in K}$, $\{g_{j}\}_{j\in J}$ are arbitrary orthonormal bases of $\mathcal{Z}$, $\mathcal{H}$, respectively555 The idea—which I include because I could not find a reference which covers this result for non-separable Hilbert spaces—is that the span of the (obviously orthogonal) set $\{|f_{k}\rangle\langle g_{j}|\}_{j\in J,k\in K}$ is dense in $\mathcal{B}^{2}(\mathcal{H},\mathcal{Z})$ (actually: dense in $\mathcal{B}^{p}$ for all $p\in[1,\infty]$, cf. Coro. 2 in A). Hence it is an orthonormal basis (Ringrose, 1971, Thm. 1.6.3). .
Another important concept we need in this regard is tensor product of Hilbert spaces and of operators thereon. The summary of what we need—where for details we refer to (Kadison and Ringrose, 1983, Ch. 2.6) or (vom Ende, 2020, Appendix A.3)—reads as follows: We will write $\mathcal{H}\otimes\mathcal{Z}$ for the (Hilbert space) tensor product of $\mathcal{H},\mathcal{Z}$; then, given any orthonormal bases $\{g_{j}\}_{j\in J}$, $\{f_{k}\}_{k\in K}$ of $\mathcal{H}$, $\mathcal{Z}$, respectively, $\{g_{j}\otimes f_{k}\}_{j\in J,k\in K}$ is an orthonormal basis of $\mathcal{H}\otimes\mathcal{Z}$. For bounded operators $B_{1}\in\mathcal{B}(\mathcal{H}_{1},\mathcal{Z}_{1})$, $B_{2}\in\mathcal{B}(\mathcal{H}_{2},\mathcal{Z}_{2})$ there exists unique $B_{1}\otimes B_{2}\in\mathcal{B}(\mathcal{H}_{1}\otimes\mathcal{H}_{2},\mathcal{Z}_{1}\otimes\mathcal{Z}_{2})$ such that $(B_{1}\otimes B_{2})(x_{1}\otimes x_{2})=B_{1}x_{1}\otimes B_{2}x_{2}$ for all $x_{1}\in\mathcal{H}_{1}$, $x_{2}\in\mathcal{H}_{2}$. Moreover, this tensor product of operators is bilinear and satisfies $(B_{1}\otimes B_{2})(B_{3}\otimes B_{4})=B_{1}B_{3}\otimes B_{2}B_{4}$, $(B_{1}\otimes B_{2})^{*}=B_{1}^{*}\otimes B_{2}^{*}$, and $\|B_{1}\otimes B_{2}\|_{\infty}=\|B_{1}\|_{\infty}\|B_{2}\|_{\infty}$ for suitable $B_{1},B_{2},B_{3},B_{4}$. If $B_{1},B_{2}$ are trace class, then so is $B_{1}\otimes B_{2}$ because of $\|B_{1}\otimes B_{2}\|_{1}=\|B_{1}\|_{1}\|B_{2}\|_{1}$, and it holds that $\operatorname{tr}(B_{1}\otimes B_{2})=\operatorname{tr}(B_{1})\operatorname{tr}(B_{2})$. Moreover, $\mathcal{B}^{2}(\mathcal{H}\otimes\mathcal{Z})$ is isometrically isomorphic (in the sense of Hilbert spaces) to $\mathcal{B}^{2}(\mathcal{H})\otimes\mathcal{B}^{2}(\mathcal{Z})$ for all complex Hilbert spaces $\mathcal{H},\mathcal{Z}$. This is due to the fact that $\{|f_{j}\rangle\langle f_{k}|\otimes|g_{a}\rangle\langle g_{b}|\}_{j,k\in J,a,b\in A}=\{|f_{j}\otimes g_{a}\rangle\langle f_{k}\otimes g_{b}|\}_{j,k\in J,a,b\in A}$ is an orthonormal basis for both of these spaces, where $\{f_{j}\}_{j\in J}$, $\{g_{a}\}_{a\in A}$ is any orthonormal basis of $\mathcal{H}$, $\mathcal{Z}$, respectively.
Operators on Schatten Classes
The objects central to our main results are linear maps between operator spaces (sometimes called superoperators). First, given any $\Phi\in\mathcal{B}(\mathcal{B}^{p}(\mathcal{H}),\mathcal{B}^{q}(\mathcal{Z}))$, $p,q\geq 1$ we denote the operator norm $\sup_{A\in\mathcal{B}^{p}(\mathcal{H}),\|A\|_{p}=1}\|\Phi(A)\|_{q}$ by $\|\Phi\|_{p\to q}$ ; similarly we write $\|\Phi\|_{\infty\to\infty}$ for the norm of any $\Phi\in\mathcal{B}(\mathcal{B}(\mathcal{H}),\mathcal{B}(\mathcal{Z}))$. Next, some $\Phi\in\mathcal{L}(\mathcal{B}^{1}(\mathcal{H}),\mathcal{B}^{1}(\mathcal{Z}))$ is called
  • 1.
    positive if for all $A\in\mathcal{B}^{1}(\mathcal{H})$ positive semi-definite (i.e. $A$ is self-adjoint with $\langle x,Ax\rangle\geq 0$ for all $x\in\mathcal{H}$, denoted by $A\geq 0$, sometimes called “PSD”) one has $\Phi(A)\geq 0$.
  • 2.
    $n$-positive for some $n\in\mathbb{N}$, if ${\rm id}_{n}\otimes\Phi:\mathcal{B}^{1}(\mathbb{C}^{n}\otimes\mathcal{H})\to\mathcal{B}^{1}(\mathbb{C}^{n}\otimes\mathcal{Z})$ defined via666 Here one uses implicitly that $\mathbb{C}^{n}\otimes\mathcal{H}\simeq\mathcal{H}\times\ldots\times\mathcal{H}$ (Kadison and Ringrose, 1983, Rem. 2.6.8) so $\mathcal{B}(\mathbb{C}^{n}\otimes\mathcal{H})$ can be identified with $\mathbb{C}^{n\times n}\otimes\mathcal{B}(\mathcal{H})$ (Kadison and Ringrose, 1983, p. 147 ff.), and similarly for the trace class.
    $({\rm id}_{n}\otimes\Phi)\begin{pmatrix}A_{11}&\cdots&A_{1n}\\ \vdots&\ddots&\vdots\\ A_{n1}&\cdots&A_{nn}\end{pmatrix}:=\begin{pmatrix}\Phi(A_{11})&\cdots&\Phi(A_{1n})\\ \vdots&\ddots&\vdots\\ \Phi(A_{n1})&\cdots&\Phi(A_{nn})\end{pmatrix}$
    for all $\{A_{jk}\}_{j,k=1}^{n}\subset\mathcal{B}^{1}(\mathcal{H})$ is positive for all $n\in\mathbb{N}$.
  • 3.
    completely positive if $\Phi$ is $n$-positive for all $n\in\mathbb{N}$. We denote the set of all completely positive maps $\Phi:\mathcal{B}^{1}(\mathcal{H})\to\mathcal{B}^{1}(\mathcal{Z})$ by $\mathsf{CP}(\mathcal{H},\mathcal{Z})$.
The notions for maps $\Phi\in\mathcal{L}(\mathcal{B}(\mathcal{H}),\mathcal{B}(\mathcal{Z}))$ are analogous by means of the isometric isomorphism $\mathcal{B}(\mathcal{Z},\mathcal{H})\simeq(\mathcal{B}^{1}(\mathcal{H},\mathcal{Z}))^{\prime}$, $B\mapsto{\rm tr}(B(\cdot))$. Indeed, this not only translates the $*{}^{*}$-topology on $(\mathcal{B}^{1}(\mathcal{H},\mathcal{Z}))^{\prime}$ into a topology on $\mathcal{B}(\mathcal{Z},\mathcal{H})$—called the ultraweak topology777 i.e. a net $\{B_{j}\}_{j\in J}\subseteq\mathcal{B}(\mathcal{Z},\mathcal{H})$ converges to $B\in\mathcal{B}(\mathcal{Z},\mathcal{H})$ in the ultraweak topology (“ultraweakly”) if and only if $\{{\rm tr}(B_{j}A)\}_{j\in J}$ converges to ${\rm tr}(BA)$ for all $A\in\mathcal{B}^{1}(\mathcal{H},\mathcal{Z})$. —but every $\Phi\in\mathcal{B}(\mathcal{B}^{1}(\mathcal{H}),\mathcal{B}^{1}(\mathcal{Z}))$ induces a unique dual map $\Phi^{*}\in\mathcal{B}(\mathcal{B}(\mathcal{Z}),\mathcal{B}(\mathcal{H}))$ via ${\rm tr}(\Phi(A)B)={\rm tr}(A\Phi^{*}(B))$ for all $A\in\mathcal{B}^{1}(\mathcal{H})$, $B\in\mathcal{B}(\mathcal{Z})$. Then $\|\Phi\|_{1\to 1}=\|\Phi^{*}\|_{\infty\to\infty}$ and (complete) positivity of $\Phi$ is well known to be equivalent to (complete) positivity of $\Phi^{*}$.
3 Recap: The Choi Matrix in Infinite Dimensions
As discussed before we need to make sense of the Choi matrix $({\rm id}\otimes\Phi)(|\Gamma\rangle\langle\Gamma|)$ in infinite dimensions, and there are different ways to go about this. Two approaches where none of the involved objects get modified are as follows: One can either consider the quadratic form induced by the Choi matrix—called Choi-Jamiołkowski form—cf. Holevo (2011a, b); Haapasalo (2021), or one can take appropriate inductive limits by considering finite truncations of $|\Gamma\rangle\langle\Gamma|$ as well as the output, cf. Friedland (2019). For our purposes, however, we can take a more naive approach where one weights the input of the “usual” Choi matrix, i.e. one replaces $|\Gamma\rangle$ by $\Gamma_{\lambda,G}:=\sum_{j\in J}\lambda_{j}^{*}g_{j}\otimes g_{j}\in\mathcal{H}\otimes\mathcal{H}$ with $\lambda\in\ell^{2}(J,\mathbb{C})$ where $G:=\{g_{j}\}_{j\in J}$ is any orthonormal basis of $\mathcal{H}$, cf. Li and Du (2015). The state $\Gamma_{\lambda,G}$ is also known in the physics literature as two-mode squeezed vacuum state, cf. Schumaker and Caves (1985). Indeed, this physical perspective on $\Gamma_{\lambda,G}$ has led to this approach being widely adopted in, e.g., quantum optics, cf. Pirandola et al. (2017); Kiukas et al. (2017) as well as Ch. 5.2 in Serafini (2017). One drawback of this weighting approach is that for general $\Phi\in\mathcal{B}(\mathcal{B}^{1}(\mathcal{H}))$ the object $({\rm id}\otimes\Phi)(|\Gamma_{\lambda,G}\rangle\langle\Gamma_{\lambda,G}|)$ may be problematic: Indeed, if $\Phi$ is the transposition map (w.r.t. an arbitrary but fixed orthonormal basis), then $\Phi$ is positive and trace-preserving but the corresponding operator $({\rm id}\otimes\Phi)(|\Gamma_{\lambda,G}\rangle\langle\Gamma_{\lambda,G}|)$ is only densely defined, cf. Tomiyama (1983); Paulsen (2003).
The first option for guaranteeing existence of $({\rm id}\otimes\Phi)(|\Gamma_{\lambda,G}\rangle\langle\Gamma_{\lambda,G}|)$ is to restrict $\Phi$ to completely bounded maps [^8] A map $\Phi\in\mathcal{B}(\mathcal{B}^{1}(\mathcal{H}))$ is called completely bounded if $\sup_{n\in\mathbb{N}}\|{\rm id}_{n}\otimes\Phi\|_{1\to 1}<\infty$., cf. Størmer (2015); Magajna (2021); Han et al. (2023). This is usually done in the framework of von Neumann algebras and factors, where the explicit vector $\Gamma_{\lambda,G}$ from above is replaced by an abstract separating and cyclic vector of the factor at hand (in our case: ${\bf 1}\otimes\mathcal{B}(\mathcal{H})$). However, while every completely positive map is completely bounded (Paulsen, 2003, Prop. 3.6), the completely bounded maps are nowhere dense in $\mathcal{B}(\mathcal{B}^{1}(\mathcal{H}))$ (Smith, 1983, Thm. 2.4 & 2.5). In contrast, the second way to make sense of the Choi formalism is the one of Li and Du (2015) where one defines the Choi operator via an appropriate infinite sum. This is also the route we will take as it is precisely what we will need when proving our main result in Section 5. Because we are allowing for non-separable Hilbert spaces we also provide a sketch of the proof for the following lemma.
Lemma 2.
Given any complex Hilbert spaces $\mathcal{H},\mathcal{Z}$, any orthonormal basis $G:=\{g_{j}\}_{j\in J}$ of $\mathcal{H}$, as well as any $\lambda\in\ell^{2}(J,\mathbb{C})$ the map
$\mathsf{C}_{\lambda,G}:\mathcal{B}(\mathcal{B}^{1}(\mathcal{H}),\mathcal{B}^{1}(\mathcal{Z}))$ $\to\mathcal{B}^{2}(\mathcal{H}\otimes\mathcal{Z})$
$\Phi$ $\mapsto\sum_{j,k\in J}\lambda_{j}^{*}\lambda_{k}|g_{j}\rangle\langle g_{k}|\otimes\Phi(|g_{j}\rangle\langle g_{k}|)$
is well-defined, linear, and bounded with $\|\mathsf{C}_{\lambda,G}\|\leq\|\lambda\|_{2}^{2}$. Moreover, $\mathsf{C}_{\lambda,G}(\Phi)$ is positive semi-definite whenever $\Phi\in\mathsf{CP}(\mathcal{H},\mathcal{Z})$.
Proof.
The key observation is that $\{\lambda_{j}^{*}\lambda_{k}|g_{j}\rangle\langle g_{k}|\otimes\Phi(|g_{j}\rangle\langle g_{k}|):(j,k)\in J\times J\}$ is an orthogonal subset of $\mathcal{B}^{2}(\mathcal{H}\otimes\mathcal{Z})$ due to $|g_{j}\rangle\langle g_{k}|\in\mathcal{B}^{1}(\mathcal{H})\subseteq\mathcal{B}^{2}(\mathcal{H})$ and $\Phi(|g_{j}\rangle\langle g_{k}|)\in\mathcal{B}^{1}(\mathcal{Z})\subseteq\mathcal{B}^{2}(\mathcal{Z})$. Thus by (Ringrose, 1971, Lemma 1.6.1) $\{\lambda_{j}^{*}\lambda_{k}|g_{j}\rangle\langle g_{k}|\otimes\Phi(|g_{j}\rangle\langle g_{k}|:j,k\in J\}$ is summable in $\mathcal{B}^{2}(\mathcal{H}\otimes\mathcal{Z})$ if and only if $\sum_{j,k\in J}\|\lambda_{j}^{*}\lambda_{k}|g_{j}\rangle\langle g_{k}|\otimes\Phi(|g_{j}\rangle\langle g_{k}|)\|_{2}^{2}<\infty$. But—again using Lemma 1 (iii)—the latter sum evaluates to
$\sum_{j,k\in J}\|\lambda_{j}^{*}\lambda_{k}|g_{j}\rangle\langle g_{k}|\otimes\Phi(|g_{j}\rangle\langle g_{k}|)\|_{2}^{2}$ $=\sum_{j,k\in J}|\lambda_{j}|^{2}|\lambda_{k}|^{2}\,\big{\|}|g_{j}\rangle\langle g_{k}|\big{\|}_{2}^{2}\big{\|}\Phi(|g_{j}\rangle\langle g_{k}|)\big{\|}_{2}^{2}$
$\leq\sum_{j,k\in J}|\lambda_{j}|^{2}|\lambda_{k}|^{2}\big{\|}\Phi(|g_{j}\rangle\langle g_{k}|)\big{\|}_{1}^{2}$
$\leq\sum_{j,k\in J}|\lambda_{j}|^{2}|\lambda_{k}|^{2}\|\Phi\|_{1\to 1}^{2}\big{\|}|g_{j}\rangle\langle g_{k}|\big{\|}_{1}^{2}$
$=\|\Phi\|_{1\to 1}^{2}\|\lambda\|_{2}^{4}<\infty\,.$
Hence $\mathsf{C}_{\lambda,G}$ is well defined and linear, and it is—again by (Ringrose, 1971, Lemma 1.6.1)—bounded with $\|\mathsf{C}_{\lambda,G}(\Phi)\|_{2}^{2}\leq\|\Phi\|_{1\to 1}^{2}\|\lambda\|_{2}^{4}$, i.e. $\|\mathsf{C}_{\lambda,G}\|\leq\|\lambda\|_{2}^{2}$.
For the final statement, given any $F\subseteq J$ non-empty and finite define $\psi_{F}:=\sum_{j\in F}\lambda_{j}^{*}|j\rangle\otimes g_{j}\in\mathbb{C}^{|F|}\otimes\mathcal{H}$, and define $U_{F}:{\rm span}\{g_{j}:j\in F\}\to\mathbb{C}^{|F|}$ to be the unique linear map such that $Ug_{j}=|j\rangle$ for all $j\in F$, i.e. $U_{F}$ is unitary. With this, one readily verifies that $\sum_{j,k\in F}\lambda_{j}^{*}\lambda_{k}|g_{j}\rangle\langle g_{k}|\otimes\Phi(|g_{j}\rangle\langle g_{k}|)$ is equal to $(U_{F}\otimes{\bf 1}_{\mathcal{H}})^{*}(({\rm id}_{|F|}\otimes\Phi)(|\psi_{F}\rangle\langle\psi_{F}|))(U_{F}\otimes{\bf 1}_{\mathcal{H}})$. But the latter is positive semi-definite because $\Phi$ is completely positive; thus the same holds for the (strong and thus weak) limit $\mathsf{C}_{\lambda,G}(\Phi)$ of $\{\sum_{j,k\in F}\lambda_{j}^{*}\lambda_{k}|g_{j}\rangle\langle g_{k}|\otimes\Phi(|g_{j}\rangle\langle g_{k}|)\}_{F\subseteq J\text{ finite}}$, as follows from the fact that any $X\in\mathcal{B}^{2}(\mathcal{H})$ is positive semi-definite if and only if ${\rm tr}(BX)\geq 0$ for all $B\in\mathcal{B}^{2}(\mathcal{H})$, $B\geq 0$. ∎
We waived the converse of this lemma (i.e. Choi operator being positive implies complete positivity, provided $\lambda_{j}\neq 0$ for all $j\in J$) because we will not need it in this work, and because $\lambda\in\ell^{2}(J,\mathbb{C}\setminus\{0\})$ only makes sense in the separable case anyway which is already covered, e.g., by (Li and Du, 2015, Thm. 1.4).
Now one feature of the Choi formalism in finite dimensions—which is important in quantum information theory— is that it establishes a one-to-one relation between completely positive and trace-preserving linear maps—also known as “CPTP maps” or “quantum channels”—and quantum states (positive semi-definite trace-class operators of unit trace) on the larger space $\mathcal{H}\otimes\mathcal{H}$ which satisfy999 In what follows, given any $A\in\mathcal{B}^{1}(\mathcal{H}\otimes\mathcal{Z})$ we write ${\rm tr}_{\mathcal{H}}(A)$ for the unique operator in $\mathcal{B}^{1}(\mathcal{Z})$ which satisfies ${\rm tr}({\rm tr}_{\mathcal{H}}(A)B)={\rm tr}(A({\bf 1}_{\mathcal{H}}\otimes B))$ for all $B\in\mathcal{B}(\mathcal{Z})$. ${\rm tr}_{\mathcal{H}}(\rho)=\frac{{\bf 1}_{\mathcal{Z}}}{\dim(\mathcal{Z})}$ (Heinosaari and Ziman, 2012, Thm. 4.48). More generally, the one-to-one correspondence in finite dimensions is between completely positive maps and positive semi-definite matrices, assuming $\lambda_{j}\neq 0$ for all $j$. There are two possible approaches of extending this to infinite dimensions: either one restricts the domain of $\mathsf{C}_{\lambda,G}$ to the completely bounded maps, or one restricts the sequence $\lambda$ to something absolutely summable. Indeed, the previously discussed transposition map example shows that $\lambda\in\ell^{1}(J,\mathbb{C})$ is the “best possible choice”: for no $\lambda\in\ell^{p}(J,\mathbb{C})\setminus\ell^{1}(J,\mathbb{C})$ with $p>1$ arbitrary would $\mathsf{C}_{\lambda,G}$ (with co-domain $\mathcal{B}^{1}$) be well defined. As the path via completely bounded maps has been sufficiently explored already (more on this in a bit) we will pursue the $\ell^{1}$-approach. Doing so—like in Lemma 2—yields a well-defined map
$\begin{split}\mathsf{C}_{\lambda,G}:\mathcal{B}(\mathcal{B}^{1}(\mathcal{H}),\mathcal{B}^{1}(\mathcal{Z}))&\to\mathcal{B}^{1}(\mathcal{H}\otimes\mathcal{Z})\\ \Phi&\mapsto\sum_{j,k\in J}\lambda_{j}^{*}\lambda_{k}|g_{j}\rangle\langle g_{k}|\otimes\Phi(|g_{j}\rangle\langle g_{k}|)\,.\end{split}$ (1)
However, even with this modification in place it turns out that the channel-state (and even the CP-PSD) duality of the Choi formalism is a purely finite-dimensional effect. More precisely, injectivity of $\mathsf{C}_{\lambda,G}$ needs separability of $\mathcal{H}$, and surjectivity never holds as soon as $\mathcal{H}$ is infinite-dimensional. This complements recent, similar results for the completely bounded approach (Han et al., 2023, Thms. 2.2 & 3.3), and this will be an important insight when discussing possible generalizations of our main result later on (cf. Sec. 6).
Proposition 1.
Given any complex Hilbert spaces $\mathcal{H},\mathcal{Z}$, any orthonormal basis $G:=\{g_{j}\}_{j\in J}$ of $\mathcal{H}$, and any $\lambda\in\ell^{1}(J,\mathbb{C})$ the map $\mathsf{C}_{\lambda,G}$ from Eq. (1) is
  • (i)
    injective if and only if $\mathcal{H}$ is separable and $\lambda_{j}\neq 0$ for all $j\in J$.
  • (ii)
    surjective if and only if $\mathcal{H}$ is finite-dimensional and $\lambda_{j}\neq 0$ for all $j\in J$.
Moreover if $\dim\mathcal{H}=\infty$, then there exist positive semi-definite trace-class operators on $\mathcal{H}\otimes\mathcal{Z}$ which have empty pre-image under $\mathsf{C}_{\lambda,G}$, regardless of the chosen $\lambda\in\ell^{1}(J,\mathbb{C})$ and the chosen orthonormal basis $\{g_{j}\}_{j\in J}$ of $\mathcal{H}$.
Proof.
(i): $\Rightarrow$”: If $\mathcal{H}$ is not separable, then $J$ is uncountable meaning there has to exist $j\in J$ such that $\lambda_{j}=0$ (Ringrose, 1971, Lemma 1.2.7). Thus it suffices to show that $\mathsf{C}_{\lambda,G}$ has non-trivial kernel as soon as $\lambda_{j}=0$ for some $j\in J$. For this note that $\Phi_{j}\in\mathcal{B}(\mathcal{B}^{1}(\mathcal{H}),\mathcal{B}^{1}(\mathcal{Z}))$ defined via $\Phi_{j}(X):=\langle g_{j},Xg_{j}\rangle Z$ (where $Z\in\mathcal{B}^{1}(\mathcal{Z})\setminus\{0\}$ is arbitrary but fixed, so $\Phi_{j}\neq 0$) is in the kernel of $\mathsf{C}_{\lambda,G}$ due to $\mathsf{C}_{\lambda,G}(\Phi_{j})=|\lambda_{j}|^{2}|g_{j}\rangle\langle g_{j}|\otimes Z=0$.
$\Leftarrow$”: Given two elements $\Phi_{1}\neq\Phi_{2}$ from $\mathcal{B}(\mathcal{B}^{1}(\mathcal{H}),\mathcal{B}^{1}(\mathcal{Z}))$, Coro. 2 (A) implies that $\Phi_{1}(|g_{j}\rangle\langle g_{k}|)\neq\Phi_{2}(|g_{j}\rangle\langle g_{k}|)$ for some $j,k\in J$ (because $\operatorname{span}\{|g_{j}\rangle\langle g_{k}|:j,k\in J\}$ is dense in $(\mathcal{B}^{1}(\mathcal{H}),\|\cdot\|_{1})$). Hence for some $x,y\in\mathcal{Z}$
$(\lambda_{j}^{*}\lambda_{k})^{-1}\langle g_{j}\otimes x,\mathrm{C}_{\lambda}(\Phi_{1})(g_{k}\otimes y)\rangle$ $=\langle x,\Phi_{1}(|g_{j}\rangle\langle g_{k}|)y\rangle\neq\langle x,\Phi_{2}(|g_{j}\rangle\langle g_{k}|)y\rangle$
$=(\lambda_{j}^{*}\lambda_{k})^{-1}\langle g_{j}\otimes x,\mathrm{C}_{\lambda}(\Phi_{2})(g_{k}\otimes y)\rangle\,.$
Thus the assumption $\lambda_{j},\lambda_{k}\neq 0$ implies $\mathrm{C}_{\lambda}(\Phi_{1})\neq\mathrm{C}_{\lambda}(\Phi_{2})$ as desired.
(ii): $\Leftarrow$”: Let $X\in\mathcal{B}^{1}(\mathcal{H}\otimes\mathcal{Z})$. Because $\lambda_{j}\neq 0$ for all $j\in J$ and because $\dim\mathcal{H}<\infty$ one may define a unique (bounded, because finite-dimensional domain) linear map $\Phi_{X}:\mathcal{B}^{1}(\mathcal{H})=\mathbb{C}^{|J|\times|J|}\to\mathcal{B}^{1}(\mathcal{Z})$ via the relation $\Phi_{X}(|g_{j}\rangle\langle g_{k}|)=(\lambda_{j}^{*}\lambda_{k})^{-1}{\rm tr}_{\mathcal{H}}((|g_{k}\rangle\langle g_{j}|\otimes{\bf 1}_{\mathcal{Z}})X)$ for all $j,k\in J$ (also recall footnote 9). A straightforward computation then shows $\mathsf{C}_{\lambda,G}(\Phi_{X})=X$.
$\Rightarrow$”: We argue by contraposition, so we have to take care of two cases:
  • 1.
    If there exists $j\in J$ such that $\lambda_{j}=0$, then $\mathsf{C}_{\lambda,G}^{-1}(|g_{j}\rangle\langle g_{j}|\otimes Z)=\emptyset$ for arbitrary but fixed $Z\in\mathcal{B}^{1}(\mathcal{H})\setminus\{0\}$ (which shows that $\mathsf{C}_{\lambda,G}^{-1}$ is not surjective): Assume to the contrary that there exists $\Phi\in\mathcal{B}(\mathcal{B}^{1}(\mathcal{H}),\mathcal{B}^{1}(\mathcal{Z}))$ such that $\mathsf{C}_{\lambda,G}(\Phi)=|g_{j}\rangle\langle g_{j}|\otimes Z$. Given any $x,y\in\mathcal{Z}$ such that $\langle x,Zy\rangle\neq 0$—which exist because $Z\neq 0$—we compute
    $0\neq\langle x,Zy\rangle$ $=\big{\langle}g_{j}\otimes x,(|g_{j}\rangle\langle g_{j}|\otimes Z)(g_{j}\otimes y)\big{\rangle}$
    $=\big{\langle}g_{j}\otimes x,\mathsf{C}_{\lambda,G}(\Phi)(g_{j}\otimes y)\big{\rangle}=|\lambda_{j}|^{2}\langle x,\Phi(|g_{j}\rangle\langle g_{j}|)y\rangle\,,$
    contradicting $\lambda_{j}=0$.
  • 2.
    Assume $\dim{\mathcal{H}}=\infty$ and w.l.o.g. $\mathbb{N}\subseteq J$. Given any $z\in\mathcal{Z}\setminus\{0\}$ we claim that $Y:=(\sum_{j\in\mathbb{N}}\frac{1}{j^{2}}|g_{j}\rangle\langle g_{j}|)\otimes|z\rangle\langle z|\in\mathcal{B}^{1}(\mathcal{H}\otimes\mathcal{Z})$ has empty pre-image under $\mathsf{C}_{\lambda,G}$, regardless of the chosen $G,\lambda$. Assume to the contrary that there exists $\Phi\in\mathcal{B}(\mathcal{B}^{1}(\mathcal{H}),\mathcal{B}^{1}(\mathcal{Z}))$ such that $\mathsf{C}_{\lambda,G}(\Phi)=Y$. This for all $j\in\mathbb{N}$ implies
    $\|z\|^{2}j^{-2}$ $=\langle g_{j}\otimes z,Y(g_{j}\otimes z)\rangle$
    $=\langle g_{j}\otimes z,\mathsf{C}_{\lambda,G}(\Phi)(g_{j}\otimes z)\rangle=|\lambda_{j}|^{2}\langle z,\Phi(|g_{j}\rangle\langle g_{j}|)z\rangle$
    so in particular $\lambda_{j}\neq 0$ for all $j\in\mathbb{N}$. Together with boundedness of $\Phi$ this yields $\infty>\|\Phi\|_{1\to 1}\|z\|^{2}\geq\sup_{j\in\mathbb{N}}|\langle z,\Phi(|g_{j}\rangle\langle g_{j}|)z\rangle|=\sup_{j\in\mathbb{N}}\tfrac{\|z\|^{2}}{j^{2}|\lambda_{j}|^{2}}$, i.e. there exists $C>0$ such that $\frac{1}{j^{2}|\lambda_{j}|^{2}}\leq C$ for all $j\in\mathbb{N}$. Therefore $\infty=C^{-1/2}\sum_{j\in\mathbb{N}}\frac{1}{j}\leq\sum_{j\in\mathbb{N}}|\lambda_{j}|\leq\|\lambda\|_{1}$, a contradiction.
The second case in the proof of (ii) also proves the additional statement: $Y$ is positive semi-definite, but its pre-image under $\mathsf{C}_{\lambda,G}$ would formally read $X\mapsto{\rm tr}(\Lambda X)|z\rangle\langle z|$ where $\Lambda:=\sum_{j\in\mathbb{N}}(j|\lambda_{j}|)^{-2}|g_{j}\rangle\langle g_{j}|$ (this map is similar in spirit to “reset channels” from quantum information). However, $\lambda\in\ell^{1}(J,\mathbb{C})$ forces $\Lambda$ to be unbounded meaning the formal map $X\mapsto{\rm tr}(\Lambda X)|z\rangle\langle z|$ is not well defined if and only if ${\rm dim}(\mathcal{H})=\infty$. ∎
Note that lack of surjectivity does not come from our choice to go from $\lambda\in\ell^{2}(J,\mathbb{C})$ to $\ell^{1}(J,\mathbb{C})$. This can be seen by adjusting the argument from the proof of Proposition 1 to $Y=\sum_{j}|\lambda_{j}|^{2}|g_{j}\rangle\langle g_{j}|\otimes Z\in\mathcal{B}^{2}(\mathcal{H}\otimes\mathcal{Z})$ for some $Z\in\mathcal{B}^{2}(\mathcal{Z})\setminus\mathcal{B}^{1}(\mathcal{H})$ and any $\lambda\in\ell^{2}(J,\mathbb{C})$: The idea is that the pre-image of $Y$ under $\mathsf{C}_{\lambda,G}$ would be $X\mapsto{\rm tr}(X)Z$ which is not in the domain of $\mathsf{C}_{\lambda,G}$ as $Z$ is not trace class. Even worse, as mentioned before the failure of the channel-state-duality also persists when restricting $\mathsf{C}_{\lambda,G}$ to a map between the completely bounded maps and the trace class on $\mathcal{H}\otimes\mathcal{Z}$ (Han et al., 2023, Thm. 3.3).
4 (Super)operators on Schatten Classes and Their Operator-Sum Forms
The key feature of the Choi formalism is that one can explicitly construct so-called Kraus operators of any completely positive linear map from it, as first shown by Choi (1975). More precisely, given any $\Phi\in\mathsf{CP}(\mathbb{C}^{n},\mathbb{C}^{k})$ the term “Kraus operators” refers to any finite family $\{V_{j}\}_{j\in J}\subset\mathbb{C}^{k\times n}$ such that $\Phi=\sum_{j\in J}V_{j}(\cdot)V_{j}^{*}$, cf. (Heinosaari and Ziman, 2012, Ch. 4.2). Interestingly, this characterization is well known to carry over to infinite dimensions. However, there one has to be careful about how to interpret the—possibly uncountably infinite—sum $\sum_{j\in J}V_{j}(\cdot)V_{j}^{*}$. This is where common topologies on $\mathcal{B}(\mathcal{X},\mathcal{Y})$ (with $\mathcal{X},\mathcal{Y}$ normed spaces) weaker than the norm topology come into play, more precisely the strong operator topology—which is the topology induced by the seminorms $\{T\mapsto\|Tx\|\}_{x\in\mathcal{X}}$—and the weak operator topology—which is induced by the seminorms $\{T\mapsto|\varphi(Tx)|\}_{x\in\mathcal{X},\varphi\in\mathcal{Y}^{*}}$ (where $\mathcal{Y}^{*}$ denotes the topological dual of $\mathcal{Y}$). If $\mathcal{Y}$ is a Hilbert space the seminorms inducing the weak operator topology reduce to $|\langle y,Tx\rangle|$. For more details we refer, e.g., to (Dunford and Schwartz, 1958, Ch. VI.1) or (vom Ende, 2020, Prop. 2.1.20).
The first technicality regarding the Kraus form—sometimes also called operator-sum form—in infinite dimensions is that for maps $\Phi\in\mathsf{CP}(\mathcal{H},\mathcal{Z})$ which are also trace-preserving the Kraus operators satisfy $\sum_{j\in J}V_{j}^{*}V_{j}={\bf 1}_{\mathcal{H}}$ in addition. Because we allow for non-separable spaces we choose to be explicit about how such sums are to be understood. The statement here—which goes back to Vigier (1946) and can be found in various versions in (Davies, 1976, Ch. 1.6), (Sakai, 1971, Lemma 1.7.4), (Kato, 1980, Ch. 8, Thm. 3.3), and (Dixmier, 1981, Appendix II)—is that norm-bounded increasing nets of self-adjoint operators converge automatically in various topologies.
Lemma 3.
Let complex Hilbert spaces $\mathcal{H},\mathcal{Z}$ as well as $\{V_{j}\}_{j\in J}\subset\mathcal{B}(\mathcal{H},\mathcal{Z})$ be given. Assume that $\{\sum_{j\in F}V_{j}^{*}V_{j}\}_{F\subseteq J\text{ finite}}$ is uniformly bounded, i.e. there exists $C>0$ such that $\|\sum_{j\in F}V_{j}^{*}V_{j}\|_{\infty}\leq C$ for all $F\subseteq J$ finite. Then $\{\sum_{j\in F}V_{j}^{*}V_{j}\}_{F\subseteq J\text{ finite}}$ admits a supremum $X\in\mathcal{B}(\mathcal{H})$, i.e. this $X$ satisfies $\sum_{j\in F}V_{j}^{*}V_{j}\leq X$ for all $F\subseteq J$ finite, and if some $Y\in\mathcal{B}(\mathcal{H})$ satisfies $\sum_{j\in F}V_{j}^{*}V_{j}\leq Y$ for all $F\subseteq J$ finite, then $X\leq Y$. Moreover, $\{\sum_{j\in F}V_{j}^{*}V_{j}\}_{F\subseteq J\text{ finite}}$ converges to $X$ in the strong operator topology as well as the ultraweak topology. In particular, $X\geq 0$.
Proof.
Let any $F\subseteq J$ finite be given. Because $\sum_{j\in F}V_{j}^{*}V_{j}$ is self-adjoint, $\|\sum_{j\in F}V_{j}^{*}V_{j}\|_{\infty}\leq C$ is equivalent to $\sum_{j\in F}V_{j}^{*}V_{j}\leq C\cdot{\bf 1}$. With this, (Dixmier, 1981, Appendix II) guarantees the existence of the supremum $X$, and the proof of said result shows that $X$ is precisely the limit of $\{\sum_{j\in F}V_{j}^{*}V_{j}\}_{F\subseteq J\text{ finite}}$ in the strong operator topology (as $\{\tfrac{1}{C}\sum_{j\in F}V_{j}^{*}V_{j}\}_{F\subseteq J\text{ finite}}\subset N$, in Dixmier’s notation). Finally, $X\geq 0$ is a direct consequence of the supremum’s property, and ultraweak convergence is due to (Sakai, 1971, Lemma 1.7.4). ∎
In what follows, for uniformly bounded nets $\{\sum_{j\in F}V_{j}^{*}V_{j}\}_{F\subseteq J\text{ finite}}$ we denote the strong limit by $\sum_{j\in J}V_{j}^{*}V_{j}$. This condition of uniform boundedness is the key to establishing in which sense the Kraus form converges.
Lemma 4.
Let complex Hilbert spaces $\mathcal{H},\mathcal{Z}$ as well as any $\{V_{j}\}_{j\in J}\subset\mathcal{B}(\mathcal{H},\mathcal{Z})$ be given such that the set $\{\sum_{j\in F}V_{j}^{*}V_{j}\}_{F\subseteq J\text{ finite}}$ is uniformly bounded. Then the following statements hold.
  • (i)
    $\{\sum_{j\in F}V_{j}^{*}BV_{j}\}_{F\subseteq J\text{ finite}}$ converges strongly and ultraweakly for all $B\in\mathcal{B}(\mathcal{Z})$. The linear map $\Phi_{V}^{*}:\mathcal{B}(\mathcal{Z})\to\mathcal{B}(\mathcal{H})$, $B\mapsto\sum_{j\in J}V_{j}BV_{j}^{*}$ is completely positive and bounded with norm $\|\Phi_{V}^{*}\|_{\infty\to\infty}=\|\sum_{j\in J}V_{j}^{*}V_{j}\|_{\infty}$.
  • (ii)
    $\{\sum_{j\in F}V_{j}AV_{j}^{*}\}_{F\subseteq J\text{ finite}}$ converges in trace norm for all $A\in\mathcal{B}^{1}(\mathcal{H})$. The linear map $\Phi_{V}:\mathcal{B}^{1}(\mathcal{H})\to\mathcal{B}^{1}(\mathcal{Z})$, $A\mapsto\sum_{j\in J}V_{j}AV_{j}^{*}$ is completely positive and bounded with norm $\|\Phi_{V}\|_{1\to 1}=\|\sum_{j\in J}V_{j}^{*}V_{j}\|_{\infty}$.
Proof.
(i): If $B$ is self-adjoint, then $V_{j}^{*}BV_{j}\leq\|B\|_{\infty}V_{j}^{*}V_{j}$ for all $j\in J$: using Cauchy-Schwarz, one for all $x\in\mathcal{H}$ has
$\langle x,(\|B\|_{\infty}V_{j}^{*}V_{j}-V_{j}^{*}BV_{j})x\rangle$ $\geq\|B\|_{\infty}\|V_{j}x\|^{2}-\|V_{j}x\|\|BV_{j}x\|$
$\geq\|B\|_{\infty}\|V_{j}x\|^{2}-\|B\|_{\infty}\|V_{j}x\|^{2}=0\,.$
Thus, because $\{\sum_{j\in F}V_{j}^{*}V_{j}\}_{F\subseteq J\text{ finite}}$ is assumed to be uniformly bounded the same is true for $\{\sum_{j\in F}V_{j}^{*}BV_{j}\}_{F\subseteq J\text{ finite}}=\{\sum_{j\in F}(\sqrt{B}V_{j})^{*}(\sqrt{B}V_{j})\}_{F\subseteq J\text{ finite}}$ and all $B\in\mathcal{B}(\mathcal{Z})$, $B\geq 0$. Hence Lemma 3 establishes strong and ultraweak convergence to a bounded, positive semi-definite operator $\Phi_{V}^{*}(B)$. The general case follows from the well-known fact that every bounded operator can be written as the linear combination of four positive semi-definite operators (Kadison and Ringrose, 1983, Coro. 4.2.4)101010 While the cited result deals with bounded operators, from the proof it is obvious that if the operator to be decomposed is trace class, then the four decomposing operators are trace class, as well. . This defines a positive linear map $\Phi_{V}^{*}:\mathcal{B}(\mathcal{Z})\to\mathcal{B}(\mathcal{H})$, which by the Russo-Dye theorem (Paulsen, 2003, Coro. 2.9) is bounded with $\|\Phi_{V}^{*}\|_{\infty\to\infty}=\|\Phi_{V}^{*}({\bf 1})\|_{\infty}=\|\sum_{j\in J}V_{j}^{*}V_{j}\|_{\infty}$. Repeating this procedure for $\{\sum_{j\in F}({\bf 1}_{n}\otimes V_{j})^{*}B({\bf 1}_{n}\otimes V_{j})\}_{F\subseteq J\text{ finite}}$ where $B\in\mathcal{B}(\mathbb{C}^{n}\otimes\mathcal{Z})$, $n\in\mathbb{N}$ are arbitrary even proves complete positivity of $\Phi_{V}^{*}$.
(ii): As before, one first shows summability for $A\geq 0$; given this step is a bit more technical we outsourced it to A (as Lemma 10). The general case follows, again, from the decomposition into four positive semi-definite operators (cf. footnote 10) which establishes a positive—hence bounded, cf. (Li and Du, 2015, Lemma 2.3) or (vom Ende, 2020, Lemma 2.3.6)—linear operator $\Phi_{V}$. Using ultraweak convergence, we see that the map $\Phi_{V}^{*}$ from (i) is precisely the dual of $\Phi_{V}$:
${\rm tr}(B\Phi_{V}(A))=\lim_{F}{\rm tr}\Big{(}\sum_{j\in F}BV_{j}^{*}AV_{j}\Big{)}=\lim_{F}{\rm tr}\Big{(}\sum_{j\in F}V_{j}BV_{j}^{*}A\Big{)}={\rm tr}(\Phi_{V}^{*}(B)A)$
for all $A\in\mathcal{B}^{1}(\mathcal{H})$, $B\in\mathcal{B}(\mathcal{Z})$. Hence $\Phi_{V}$ is completely positive because $\Phi_{V}^{*}$ is, and $\|\Phi_{V}\|_{1\to 1}=\|\Phi_{V}^{*}\|_{\infty\to\infty}=\|\sum_{j\in J}V_{j}^{*}V_{j}\|_{\infty}$. ∎
With this we can now state how complete positivity and the Kraus form are related in general, cf. (Davies, 1976, Ch. 9.2), (Li and Du, 2015, Thm. 1.4), or (vom Ende, 2020, Prop. 2.3.10 ff.)—which will be another key ingredient in the proof of our main result: For every $\Phi\in\mathsf{CP}(\mathcal{H},\mathcal{Z})$ (resp. $\Phi\in\mathcal{L}(\mathcal{B}(\mathcal{H}),\mathcal{B}(\mathcal{Z}))$ completely positive and ultraweakly continuous) there exist $\{V_{j}\}_{j\in J}\subset\mathcal{B}(\mathcal{H},\mathcal{Z})$ such that the set $\{\sum_{j\in F}V_{j}^{*}V_{j}\}_{F\subseteq J\text{ finite}}$ is uniformly bounded, and one has $\Phi(A)=\sum_{j\in J}V_{j}AV_{j}^{*}$ for all $A\in\mathcal{B}^{1}(\mathcal{H})$ (resp. $\Phi(B)=\sum_{j\in J}V_{j}^{*}BV_{j}$ for all $B\in\mathcal{B}(\mathcal{Z})$) where the respective sums converge as described in Lemma 4. Moreover, if $\mathcal{H},\mathcal{Z}$ are both separable, then one can choose the index set $J$ to be countable. We remark that the Kraus form has also been extended to factors on separable Hilbert spaces (Han et al., 2023, Thm. 3.1).
The final thing we have to do before coming to our main result is to go one level higher as we need to make sense of the trace of superoperators $\Phi(B^{*}(\cdot)B)$, respectively establish criteria under which such a trace exists. Because $\mathcal{B}^{2}(\mathcal{H})$ is a Hilbert space, the trace class $\mathcal{B}^{1}(\mathcal{B}^{2}(\mathcal{H}))$ on such spaces is well defined. The following lemma is about when maps $\Phi(B^{*}(\cdot)B)$ are themselves trace class:
Lemma 5.
The following statements hold.
  • (i)
    For all $p>0$ and all $X,Y\in\mathcal{B}^{p}(\mathcal{H})$ one has $X^{*}(\cdot)Y\in\mathcal{B}^{p}(\mathcal{B}^{2}(\mathcal{H}))$ with $\|X^{*}(\cdot)Y\|_{p}=\|X\|_{p}\|Y\|_{p}$.
  • (ii)
    For all $\Phi\in\mathcal{B}(\mathcal{B}^{1}(\mathcal{H}))$, $B\in\mathcal{B}^{2}(\mathcal{H})$ one has $\Phi(B^{*}(\cdot)B)\in\mathcal{B}^{2}(\mathcal{B}^{2}(\mathcal{H}))$.
  • (iii)
    For all $\Phi\in\mathcal{B}(\mathcal{B}^{1}(\mathcal{H}))$, $B\in\mathcal{B}^{1}(\mathcal{H})$ one has $\Phi(B^{*}(\cdot)B)\in\mathcal{B}^{1}(\mathcal{B}^{2}(\mathcal{H}))$.
  • (iv)
    For all $K_{1},K_{2}\in\mathcal{B}(\mathcal{H})$, $B\in\mathcal{B}^{1}(\mathcal{H})$ the map $X\mapsto K_{1}B^{*}XBK_{2}$ is in $\mathcal{B}^{1}(\mathcal{B}^{2}(\mathcal{H}))$, and its trace is given by ${\rm tr}(K_{1}B^{*}){\rm tr}(BK_{2})$.
Proof.
(i): Our goal is to show that the—by Lemma 1 (iv) well-defined and bounded—linear map $X^{*}(\cdot)Y:\mathcal{B}^{2}(\mathcal{H})\to\mathcal{B}^{2}(\mathcal{H})$ is compact, and that its singular values are given by the pairwise product of the singular values of $X$ and $Y$. Given any Schmidt decompositions $X=\sum_{j\in N_{1}}s_{j}(X)|f_{j}\rangle\langle g_{j}|$, $Y=\sum_{j\in N_{2}}s_{j}(Y)|y_{j}\rangle\langle z_{j}|$, we write $|\,|g_{j}\rangle\langle z_{k}|\,\rangle\langle\,|f_{j}\rangle\langle y_{k}|\,|$ as short-hand for the map $B\mapsto\langle\,|f_{j}\rangle\langle y_{k}|,B\rangle_{\mathrm{HS}}|g_{j}\rangle\langle z_{k}|=\langle f_{j}|B|y_{k}\rangle|g_{j}\rangle\langle z_{k}|$ on $\mathcal{B}^{2}(\mathcal{H})$. This yields
$X^{*}(\cdot)Y$ $=\Big{(}\sum_{j\in N_{1}}s_{j}(X)|f_{j}\rangle\langle g_{j}|\Big{)}^{*}(\cdot)\Big{(}\sum_{k\in N_{2}}s_{k}(Y)|y_{k}\rangle\langle z_{k}|\Big{)}$
$=\sum_{(j,k)\in N_{1}\times N_{2}}s_{j}(X)s_{k}(Y)\big{|}\,|g_{j}\rangle\langle z_{k}|\,\big{\rangle}\big{\langle}\,|f_{j}\rangle\langle y_{k}|\,\big{|}\,,$
where, in particular, the latter is a Schmidt decomposition of $X^{*}(\cdot)Y$ because $\{\,|g_{j}\rangle\langle z_{k}|\,\}_{(j,k)\in N_{1}\times N_{2}}$, $\{\,|f_{j}\rangle\langle y_{k}|\,\}_{(j,k)\in N_{1}\times N_{2}}$ are orthonormal sets in $\mathcal{B}^{2}(\mathcal{H})$. Thus the singular values of $X^{*}(\cdot)Y$ are $\{s_{j}(X)s_{k}(Y)\}_{(j,k)\in N_{1}\times N_{2}}$, meaning $\|X^{*}(\cdot)Y\|_{p}=(\sum_{j\in N_{1}}(s_{j}(X))^{p})^{1/p}(\sum_{k\in N_{2}}(s_{k}(Y))^{p})^{1/p}=\|X\|_{p}\|Y\|_{p}$.
(ii): First we show that $\Phi(B^{*}(\cdot)B)$ is a bounded operator on $\mathcal{B}^{2}(\mathcal{H})$: Using Lemma 1 we upper bound $\sup_{X\in\mathcal{B}^{2}(\mathcal{H}),\|X\|_{2}\leq 1}\|\Phi(B^{*}XB)\|_{2}$ by
$\sup_{X\in\mathcal{B}^{2}(\mathcal{H}),\|X\|_{2}\leq 1}\|\Phi(B^{*}XB)\|_{1}$
$\quad\leq\|\Phi\|_{1\to 1}\sup_{X\in\mathcal{B}^{2}(\mathcal{H}),\|X\|_{2}\leq 1}\|B^{*}XB\|_{1}\leq\|\Phi\|_{1\to 1}\sup_{X\in\mathcal{B}^{2}(\mathcal{H}),\|X\|_{2}\leq 1}\|B^{*}X\|_{2}\|B\|_{2}$
$\quad\leq\|\Phi\|_{1\to 1}\sup_{X\in\mathcal{B}^{2}(\mathcal{H}),\|X\|_{2}\leq 1}\|B^{*}\|\|X\|_{2}\|B\|_{2}=\|\Phi\|_{1\to 1}\|B\|\|B\|_{2}<\infty\,.$
Next consider any Schmidt decomposition $B=\sum_{j\in N}s_{j}(B)|f_{j}\rangle\langle g_{j}|$. Completing the orthonormal system $\{f_{j}\}_{j\in N}$ to an orthonormal basis $\{f_{j}\}_{j\in J}$ of $\mathcal{H}$ gives rise to an orthonormal basis $\{|f_{k}\rangle\langle f_{j}|\}_{j,k\in J}$ of $\mathcal{B}^{2}(\mathcal{H})$. By Theorem 2.4.3 in Ringrose (1971), $\Phi(B^{*}(\cdot)B)\in\mathcal{B}^{2}(\mathcal{B}^{2}(\mathcal{H}))$ is now equivalent to $\sum_{j,k\in J}\|\Phi(B^{*}|f_{k}\rangle\langle f_{j}|B)\|_{2}^{2}<\infty$. But the latter sum evaluates to
$\sum_{j,k\in J}\|\Phi(B^{*}|f_{k}\rangle\langle f_{j}|B)\|_{2}^{2}$ $=\sum_{j,k\in N}(s_{j}(B))^{2}(s_{k}(B))^{2}\|\Phi(|g_{k}\rangle\langle g_{j}|)\|_{2}^{2}$
$\leq\|\Phi\|_{1\to 1}^{2}\Big{(}\sum_{j\in N}(s_{j}(B))^{2}\Big{)}^{2}=\|\Phi\|_{1\to 1}^{2}\|B\|_{2}^{4}<\infty$
due to $\|\cdot\|_{2}\leq\|\cdot\|_{1}$.
(iii): By Lemma 1 (ii) there exist $B_{1},B_{2}\in\mathcal{B}^{2}(\mathcal{H})$ such that $B=B_{1}B_{2}$, so we decompose $\Phi(B^{*}(\cdot)B)=\Phi((B_{1}B_{2})^{*}(\cdot)B_{1}B_{2})=\Phi(B_{2}^{*}(\cdot)B_{2})\circ(B_{1}^{*}(\cdot)B_{1})$. By (i), $B_{1}^{*}(\cdot)B_{1}\in\mathcal{B}^{2}(\mathcal{B}^{2}(\mathcal{H}))$, and the same holds for $\Phi(B_{2}^{*}(\cdot)B_{2})$ by (ii). In particular, Lemma 1 (i) implies that their composition is in $\mathcal{B}^{1}(\mathcal{B}^{2}(\mathcal{H}))$.
(iv): As $BK_{1}^{*},BK_{2}\in\mathcal{B}^{1}(\mathcal{H})$ we know $K_{1}B^{*}(\cdot)BK_{2}\in\mathcal{B}^{1}(\mathcal{B}^{2}(\mathcal{H}))$ by (i). Given any orthonormal basis $\{g_{j}\}_{j\in J}$ of $\mathcal{H}$ (so $\{|g_{k}\rangle\langle g_{j}|\}_{j,k\in J}$ is an orthonormal basis of $\mathcal{B}^{2}(\mathcal{H})$) we can evaluate the trace in question explicitly:
${\rm tr}(K_{1}B^{*}(\cdot)BK_{2})$ $=\sum_{j,k\in J}\big{\langle}\,|g_{k}\rangle\langle g_{j}|,K_{1}B^{*}|g_{k}\rangle\langle g_{j}|BK_{2}\big{\rangle}_{\sf HS}$
$=\sum_{j,k\in J}\langle g_{k},K_{1}B^{*}g_{k}\rangle\langle g_{j},BK_{2}g_{j}\rangle={\rm tr}(K_{1}B^{*}){\rm tr}(BK_{2})$
A fact that will become important later is that given $\Phi$ completely positive, the trace of $\Phi(B^{*}(\cdot)B)$ can be evaluated explicitly using the Kraus operators of $\Phi$.
Lemma 6.
Let a complex Hilbert space $\mathcal{H}$, some $B\in\mathcal{B}^{1}(\mathcal{H})$, as well as $\Phi\in\mathsf{CP}(\mathcal{H})$ be given. If $\{V_{j}\}_{j\in J}\subset\mathcal{B}(\mathcal{H})$ is any set of Kraus operators of $\Phi$, then $v:=\{{\rm tr}(B^{*}V_{j})\}_{j\in J}\in\ell^{2}(J,\mathbb{C})$. Moreover, the trace of $\Phi(B^{*}(\cdot)B)$ is equal to $\|v\|_{2}^{2}$.
Proof.
Let any $F\subseteq J$, $F\neq\emptyset$ finite as well as any Schmidt decomposition $\sum_{j\in N}s_{j}|f_{j}\rangle\langle g_{j}|$ of $B$ be given. As done before we decompose $B$ into Hilbert-Schmidt operators $B_{1}:=\sum_{j\in N}\sqrt{s_{j}}|f_{j}\rangle\langle g_{j}|$, $B_{2}:=\sum_{j\in N}\sqrt{s_{j}}|g_{j}\rangle\langle g_{j}|$, that is, $B=B_{1}B_{2}$. Using the Cauchy-Schwarz inequality on $\mathcal{B}^{2}(\mathcal{H})$ we compute
$\sum_{j\in F}|{\rm tr}(B^{*}V_{j})|^{2}$ $=\sum_{j\in F}|\langle B_{1},V_{j}B_{2}^{*}\rangle_{\sf HS}|^{2}\leq\|B_{1}\|_{2}^{2}\sum_{j\in F}{\rm tr}(B_{2}^{*}B_{2}V_{j}^{*}V_{j})\,.$
By Lemma 3 $\{\sum_{j\in F}V_{j}^{*}V_{j}\}_{F\subseteq J\text{ finite}}$ converges ultraweakly to its supremum, which shows
$\sum_{j\in F}|{\rm tr}(B^{*}V_{j})|^{2}$ $\leq\|B_{1}\|_{2}^{2}\sum_{j\in F}{\rm tr}(B_{2}^{*}B_{2}V_{j}^{*}V_{j})\leq\|B_{1}\|_{2}^{2}\Big{\|}B_{2}^{*}B_{2}\sum_{j\in F}V_{j}^{*}V_{j}\Big{\|}_{1}$
$\leq\|B_{1}\|_{2}^{2}\|B_{2}^{*}B_{2}\|_{1}\Big{\|}\sum_{j\in F}V_{j}^{*}V_{j}\Big{\|}_{\infty}\leq\|B\|_{1}^{2}\Big{\|}\sum_{j\in J}V_{j}^{*}V_{j}\Big{\|}_{\infty}.$ (2)
This implies two things:
  • 1.
    $v=\{{\rm tr}(B^{*}V_{j})\}_{j\in J}\in\ell^{2}(J,\mathbb{C})$
  • 2.
    $w:=\{{\rm tr}(B_{2}V_{j}^{*}V_{j}B_{2}^{*})\}_{j\in J}=\{\|B_{2}^{*}V_{j}\|_{2}^{2}\}_{j\in J}\in\ell^{1}(J,\mathbb{C})$
To connect $\|v\|_{2}^{2}$ to the trace of $\Phi(B^{*}(\cdot)B)$ we complete $\{f_{j}\}_{j\in N}$ to an orthonormal basis $\{f_{a}\}_{a\in A}$ of $\mathcal{H}$ (i.e. $N\subseteq A$), which lets us evaluate
${\rm tr}(\Phi(B^{*}(\cdot)B))$ $=\sum_{a,b\in A}\big{\langle}\,|f_{a}\rangle\langle f_{b}|,\Phi\big{(}B^{*}|f_{a}\rangle\langle f_{b}|B\big{)}\big{\rangle}_{\mathrm{HS}}$
$=\sum_{a,b\in N}\sqrt{s_{a}s_{b}}\langle f_{a},\Phi(B_{2}^{*}|g_{a}\rangle\langle g_{b}|B_{2})f_{b}\rangle$
$=\sum_{a,b\in N}\sum_{j\in J}\sqrt{s_{a}s_{b}}\langle f_{a},V_{j}B_{2}^{*}g_{a}\rangle\langle g_{b},B_{2}V_{j}^{*}f_{b}\rangle\,.$
Now we want to interchange the order of summation; for this we check that $\{\sqrt{s_{a}s_{b}}\langle f_{a},V_{j}B_{2}^{*}g_{a}\rangle\langle g_{b},B_{2}V_{j}^{*}f_{b}\rangle\}_{a,b\in N,j\in J}$ is summable (Ringrose, 1971, Lemma 1.2.6). Using Cauchy-Schwarz on $\ell^{2}(N,\mathbb{C})$ we compute
$\sum_{j\in J}\sum_{a,b\in N}\big{|}\sqrt{s_{a}s_{b}}\langle f_{a}$ $,V_{j}B_{2}^{*}g_{a}\rangle\langle g_{b},B_{2}V_{j}^{*}f_{b}\rangle\big{|}$
$=\sum_{j\in J}\Big{(}\sum_{a\in N}\sqrt{s_{a}}|\langle f_{a},V_{j}B_{2}^{*}g_{a}\rangle|\Big{)}^{2}$
$\leq\sum_{j\in J}\Big{(}\sum_{a\in N}(\sqrt{s_{a}})^{2}\Big{)}\Big{(}\sum_{a\in N}|\langle f_{a},V_{j}B_{2}^{*}g_{a}\rangle|^{2}\Big{)}$
$\leq\Big{(}\sum_{a\in N}s_{a}\Big{)}\sum_{j\in J}\Big{(}\sum_{a\in A}\|V_{j}B_{2}^{*}g_{a}\|^{2}\Big{)}$
$=\|B\|_{1}\sum_{j\in J}{\rm tr}(B_{2}V_{j}^{*}V_{j}B_{2}^{*})=\|B\|_{1}\|w\|_{1}<\infty\,,$
which by (Ringrose, 1971, Lemma 1.2.5) implies summability, as desired. Thus we may interchange sums which lets us arrive at
${\rm tr}(\Phi(B^{*}(\cdot)B))$ $=\sum_{j\in J}\sum_{a,b\in N}\sqrt{s_{a}s_{b}}\langle f_{a},V_{j}B_{2}^{*}g_{a}\rangle\langle g_{b},B_{2}V_{j}^{*}f_{b}\rangle$
$=\sum_{j\in J}\sum_{a,b\in J}\langle f_{a},V_{j}B_{2}^{*}B_{1}^{*}f_{a}\rangle\langle f_{b},B_{1}B_{2}V_{j}^{*}f_{b}\rangle$
$=\sum_{j\in J}|{\rm tr}(V_{j}(B_{1}B_{2})^{*})|^{2}=\sum_{j\in J}|{\rm tr}(B^{*}V_{j})|^{2}=\|v\|_{2}^{2}\,.$
5 Unique Decompositions of Generators of Completely Positive Dynamical Semigroups
Finally, we are ready to establish our main result. For this we first single out subspaces of $\mathsf{CP}(\mathcal{H})$ for which a certain weighted trace (that exists by Lemma 5 (iii)) vanishes.
Definition 1.
Let $\mathcal{H}$ be an arbitrary complex Hilbert space and let any $B\in\mathcal{B}^{1}(\mathcal{H})$ be given. Then $\mathsf{CP}_{B}(\mathcal{H}):=\{\Phi\in\mathsf{CP}(\mathcal{H})\,:\,{\rm tr}(\Phi(B^{*}(\cdot)B))=0\}$.
By Lemma 6 $\Phi\in\mathsf{CP}_{B}(\mathcal{H})$ if and only if there exist Kraus operators $\{V_{j}\}_{j\in J}$ of $\Phi$ such that ${\rm tr}(B^{*}V_{j})=0$ for all $j\in J$ (equivalently: all sets of Kraus operators satisfy this). More importantly, this trace can also be recovered using the Choi formalism and vectorization. Let us quickly recap the latter concept: In finite dimensions, given $B\in\mathbb{C}^{m\times n}$ and any orthonormal basis $G:=\{g_{j}\}_{j=1}^{n}$ of $\mathbb{C}^{n}$ one defines ${\rm vec}_{G}(B):=\sum_{k=1}^{m}g_{j}\otimes Bg_{j}\in\mathbb{C}^{n}\otimes\mathbb{C}^{m}\simeq\mathbb{C}^{mn}$ (Horn and Johnson, 1991, Ch. 4.2 ff.). This generalizes to any complex Hilbert spaces $\mathcal{H},\mathcal{Z}$ via the map ${\rm vec}_{G}:\mathcal{B}^{2}(\mathcal{H},\mathcal{Z})\to\mathcal{H}\otimes\mathcal{Z},X\mapsto\sum_{j\in J}g_{j}\otimes Xg_{j}$ (which is well-defined and unitary) where $G:=\{g_{j}\}_{j\in J}$ is any orthonormal basis of $\mathcal{H}$, refer, e.g., to Gudder (2020). Equivalently, given any $X\in\mathcal{B}^{2}(\mathcal{H})$, ${\rm vec}_{G}(X)$ is the unique element of $\mathcal{H}\otimes\mathcal{Z}$ such that $\langle g_{j}\otimes z,{\rm vec}_{G}(X)\rangle=\langle z,Xg_{j}\rangle$ for all $j\in J$, $z\in\mathcal{Z}$.
Lemma 7.
Given any complex Hilbert space $\mathcal{H}$ and any orthonormal basis $G:=\{g_{j}\}_{j\in J}$ of $\mathcal{H}$, as well as any $\lambda\in\ell^{2}(J,\mathbb{C})$ it for all $\Phi\in\mathcal{B}(\mathcal{B}^{1}(\mathcal{H}))$ and all $X,Y\in\mathcal{B}^{2}(\mathcal{H})$ holds that
${\rm tr}\big{(}\Phi((XB)^{*}(\cdot)YB)\big{)}=\langle{\rm vec}_{G}(X),\mathsf{C}_{\lambda,G}(\Phi){\rm vec}_{G}(Y)\rangle\,,$
where $B:=\sum_{j\in J}\lambda_{j}|g_{j}\rangle\langle g_{j}|\in\mathcal{B}^{2}(\mathcal{H})$.
Proof.
We begin by expanding the right-hand side:
$\langle$ ${\rm vec}_{G}(X),\mathsf{C}_{\lambda,G}(\Phi){\rm vec}_{G}(Y)\rangle$
$\ \ =\sum_{p,j,k,q\in J}\lambda_{j}^{*}\lambda_{k}\big{\langle}g_{p}\otimes Xg_{p},(|g_{j}\rangle\langle g_{k}|\otimes\Phi(|g_{j}\rangle\langle g_{k}|))(g_{q}\otimes Yg_{q})\big{\rangle}$
$\ \ =\sum_{j,k\in J}\lambda_{j}^{*}\lambda_{k}\big{\langle}Xg_{j},\Phi(|g_{j}\rangle\langle g_{k}|)Yg_{k}\big{\rangle}=\sum_{j,k\in J}\big{\langle}g_{j},X^{*}\Phi(B^{*}|g_{j}\rangle\langle g_{k}|B)Yg_{k}\big{\rangle}\,.$
Defining $\Phi_{X^{*},Y}:=X^{*}(\cdot)Y$, $\Phi_{B^{*},B}:=B^{*}(\cdot)B$—which are both elements of $\mathcal{B}^{2}(\mathcal{B}^{2}(\mathcal{H}))$ by Lemma 5 (i)—the map $\Phi_{X^{*},Y}\circ\Phi\circ\Phi_{B^{*},B}=X^{*}\Phi(B^{*}(\cdot)B)Y$ is in $\mathcal{B}^{1}(\mathcal{B}^{2}(\mathcal{H}))$ (Lemma 1 (i) and Lemma 5 (ii)). Therefore—as $\{|g_{k}\rangle\langle g_{j}|\}_{j\in J,k\in K}$ is an orthonormal basis of $\mathcal{B}^{2}(\mathcal{H})$$\sum_{j,k\in J}\langle g_{j},X^{*}\Phi(B^{*}|g_{j}\rangle\langle g_{k}|B)Yg_{k}\rangle$ is precisely the trace of $\Phi_{X^{*},Y}\circ\Phi\circ\Phi_{B^{*},B}$. Hence
$\langle{\rm vec}_{G}(X)|\mathsf{C}_{\lambda,G}(\Phi)|{\rm vec}_{G}(Y)\rangle$ $={\rm tr}(\Phi_{X^{*},Y}\circ\Phi\circ\Phi_{B^{*},B})$
$={\rm tr}(\Phi\circ\Phi_{B^{*},B}\circ\Phi_{X^{*},Y})={\rm tr}\big{(}\Phi((XB)^{*}(\cdot)YB)\big{)}\,,$
as claimed. ∎
Because the Choi operator is positive semi-definite for completely positive inputs, the above identity places a restriction on the kernel of the Choi operator in case $\Phi$ is completely positive and has vanishing weighted trace:
Lemma 8.
Let $\mathcal{H}$ be any complex Hilbert space and let $G:=\{g_{j}\}_{j\in J}$ be any orthonormal basis of $\mathcal{H}$. Given any $\lambda\in\ell^{2}(J,\mathbb{R}_{+})$, any $X\in\mathcal{B}^{2}(\mathcal{H})$, and any $\Phi\in\mathsf{CP}_{XB}(\mathcal{H})$—where $B:=\sum_{j\in J}\lambda_{j}|g_{j}\rangle\langle g_{j}|\in\mathcal{B}^{2}(\mathcal{H})$—it holds that $\mathsf{C}_{\lambda,G}(\Phi){\rm vec}_{G}(X)=0$
Proof.
Because $\Phi$ is completely positive, $\mathsf{C}_{\lambda,G}(\Phi)$ is positive semi-definite by Lemma 2. On the other hand, ${\rm tr}(\Phi((XB)^{*}(\cdot)XB))=0$ by Lemma 7 is equivalent to $\langle{\rm vec}_{G}(X),\mathsf{C}_{\lambda,G}(\Phi){\rm vec}_{G}(X)\rangle=0$. Therefore
00 0 $=\langle{\rm vec}_{G}(X),\mathsf{C}_{\lambda,G}(\Phi){\rm vec}_{G}(X)\rangle$
$=\Big{\langle}\sqrt{\mathsf{C}_{\lambda,G}(\Phi)}{\rm vec}_{G}(X),\sqrt{\mathsf{C}_{\lambda,G}(\Phi)}{\rm vec}_{G}(X)\Big{\rangle}=\Big{\|}\sqrt{\mathsf{C}_{\lambda,G}(\Phi)}{\rm vec}_{G}(X)\Big{\|}^{2}\,.$
Altogether this shows $\mathsf{C}_{\lambda,G}(\Phi){\rm vec}_{G}(X)=\sqrt{\mathsf{C}_{\lambda,G}(\Phi)}\sqrt{\mathsf{C}_{\lambda,G}(\Phi)}{\rm vec}_{G}(X)=0$, as desired. ∎
It turns out that the kernel constraint from Lemma 8 is the reason why the $K(\cdot)+(\cdot)K^{*}$-part of any generator $L\in\mathsf{L}(\mathsf{CP}(\mathcal{H}))$ is “independent” of the completely positive part, assuming the weighted trace of the latter vanishes. This is also the point where we have to invoke separability of the underlying Hilbert space:
Proposition 2.
Let an arbitrary separable complex Hilbert space $\mathcal{H}$, any $K\in\mathcal{B}(\mathcal{H})$, and any $B\in\mathcal{B}^{1}(\mathcal{H})$ with ${\rm tr}(B)\neq 0$ be given. If $K(\cdot)+(\cdot)K^{*}\in\mathsf{CP}_{B}(\mathcal{H})-\mathsf{CP}_{B}(\mathcal{H})$, then $K=i\lambda{\bf 1}$ for some $\lambda\in\mathbb{R}$.
Proof.
We start with any Schmidt decomposition $B=\sum_{j\in N}s_{j}(B)|f_{j}\rangle\langle g_{j}|$. If $\{g_{j}\}_{j\in N}$ is not yet an orthonormal basis we can extend it to an orthonormal basis $G:=\{g_{j}\}_{j\in J}$, $N\subseteq J$ of $\mathcal{H}$; in particular $J$—and hence $J\setminus N$—are countable by assumption. With this we can define $s_{j}(B):=0$ for $j\in J\setminus N$ as well as operators $B_{1}:=\sum_{j\in J}\sqrt{s_{j}(B)}|f_{j}\rangle\langle g_{j}|$, $B_{2}:=\sum_{j\in J}\lambda_{j}|g_{j}\rangle\langle g_{j}|$ where $\lambda:J\to(0,\infty)$ is defined via
$\lambda(j):=\begin{cases}\sqrt{s_{j}(B)}&\text{if }s_{j}(B)\neq 0\\ 2^{-j}&\text{else}\\ \end{cases}\,.$
Because $J$ is countable one has $\lambda\in\ell^{2}(J,(0,\infty))$ meaning $B_{1},B_{2}\in\mathcal{B}^{2}(\mathcal{H})$ and, more importantly, $B=B_{1}B_{2}$. Now the key insight is that if $K(\cdot)+(\cdot)K^{*}=\Phi_{1}-\Phi_{2}$ for some $\Phi_{1},\Phi_{2}\in\mathsf{CP}_{B}(\mathcal{H})=\mathsf{CP}_{B_{1}B_{2}}(\mathcal{H})$, then
$\mathsf{C}_{\lambda,G}(K(\cdot)+(\cdot)K^{*}){\rm vec}_{G}(B_{1})$ $=\mathsf{C}_{\lambda,G}(\Phi_{1}-\Phi_{2}){\rm vec}_{G}(B_{1})$
$=\mathsf{C}_{\lambda,G}(\Phi_{1}){\rm vec}_{G}(B_{1})-\mathsf{C}_{\lambda,G}(\Phi_{2}){\rm vec}_{G}(B_{1})=0$
by Lemma 2 & Lemma 8. In particular, for all $j,k\in J$ it holds that
00 0 $=\langle g_{k}\otimes g_{j},\mathsf{C}_{\lambda,G}(K(\cdot)+(\cdot)K^{*}){\rm vec}_{G}(B_{1})\rangle$
$=\sum_{a,b\in J}\lambda_{a}\lambda_{b}\big{\langle}g_{k}\otimes g_{j},\big{(}|g_{a}\rangle\langle g_{b}|\otimes(K|g_{a}\rangle\langle g_{b}|+|g_{a}\rangle\langle g_{b}|K^{*})\big{)}{\rm vec}_{G}(B_{1})\big{\rangle}$
$=\sum_{b\in J}\lambda_{k}\lambda_{b}\big{(}\langle g_{j},Kg_{k}\rangle\langle g_{b},B_{1}g_{b}\rangle+\delta_{jk}\langle g_{b},K^{*}B_{1}g_{b}\rangle\big{)}\,.$
Now we use that $\lambda_{b}B_{1}g_{b}=B_{1}B_{2}g_{b}=Bg_{b}$ for all $b\in J$ to compute
00 0 $=\sum_{b\in J}\lambda_{k}\big{(}\langle g_{j},Kg_{k}\rangle\langle g_{b},Bg_{b}\rangle+\delta_{jk}\langle g_{b},K^{*}Bg_{b}\rangle\big{)}$
$=\lambda_{k}\langle g_{j},Kg_{k}\rangle\sum_{b\in J}\langle g_{b},Bg_{b}\rangle+\lambda_{k}\delta_{jk}\sum_{b\in J}\langle g_{b},K^{*}Bg_{b}\rangle$
$=\lambda_{k}\langle g_{j},Kg_{k}\rangle{\rm tr}(B)+\lambda_{k}\delta_{jk}{\rm tr}(BK^{*})\,.$ (3)
Because $\lambda_{k}\neq 0$ for all $k\in J$ by construction, (3) for all $j,k\in J$ is equivalent to $0=\langle g_{j},Kg_{k}\rangle{\rm tr}(B)+\delta_{jk}{\rm tr}(BK^{*})$. This lets us distinguish two cases: if $j\neq k$, then this computation shows $0=\langle g_{j},Kg_{k}\rangle{\rm tr}(B)$; but ${\rm tr}(B)\neq 0$ by assumption, so this even shows $\langle g_{j},Kg_{k}\rangle=0$ for all $j\neq k$. Because $\{g_{j}\}_{j\in J}$ is an orthonormal basis of $\mathcal{H}$ this means that $K$ is diagonal (w.r.t. $G$). Now if $j=k$, then (3) yields $0=\langle g_{j},Kg_{j}\rangle{\rm tr}(B)+{\rm tr}(BK^{*})$. Again, ${\rm tr}(B)\neq 0$ so we find $\langle g_{j},Kg_{j}\rangle=-\frac{{\rm tr}(BK^{*})}{{\rm tr}(B)}$; in other words there exists $c\in\mathbb{C}$ such that $\langle g_{j},Kg_{j}\rangle=c$ for all $j\in J$. Altogether, as $\{g_{j}\}_{j\in J}$ is an orthonormal basis this shows $K=c\cdot{\bf 1}$. All that is left to prove is that $c=i\lambda$ for some $\lambda\in\mathbb{R}$. Because ${\rm tr}(\Phi_{1}(B^{*}(\cdot)B))={\rm tr}(\Phi_{2}(B^{*}(\cdot)B))=0$, the same holds for $K(\cdot)+(\cdot)K^{*}$. By Lemma 5 (iv) this means $0={\rm tr}(KB^{*}){\rm tr}(B)+{\rm tr}(B^{*}){\rm tr}(BK^{*})=2{\rm Re}({\rm tr}(KB^{*}){\rm tr}(B))$. Now we insert $K=c\cdot{\bf 1}$ to find $0=2{\rm Re}(c|{\rm tr}(B)|^{2})=2|{\rm tr}(B)|^{2}{\rm Re}(c)$, and as ${\rm tr}(B)\neq 0$ this shows ${\rm Re}(c)=0$, i.e. $c=i\lambda$ for some $\lambda\in\mathbb{R}$. ∎
We now come to our main result. It turns out that the most convenient way for establishing uniqueness of decompositions $L\mapsto(K,\Phi)$ of corresponding generators is via bijectivity of a certain map (with appropriate domain and co-domain). Here, the co-domain of such a map is most easily formulated in terms of the Lie wedge which is a generalization of Lie algebras that is central to Lie semigroup theory, cf. also Hilgert et al. (1989): Given a Banach space $\mathcal{X}$ and a norm-closed (sub-)semigroup $S$ of $\mathcal{B}(\mathcal{X})$ which contains the identity, the Lie wedge of $S$ is defined as $\mathsf{L}(S):=\{A\in\mathcal{B}(\mathcal{X}):e^{tA}\in S\text{ for all }t\geq 0\}$. As an example, casting the standard forms of generators from the introduction into this language yields
$\mathsf{L}(\mathsf{CPTP}(\mathcal{H}))$ =$=\{-i[H,\cdot]+\Phi-\frac{1}{2}\{\Phi^{*}({\bf 1}),\cdot\}:H\in i\mathfrak{u}(\mathcal{H}),\Phi\in\mathsf{CP}(\mathcal{H})\}$
$\mathsf{L}(\mathsf{CP}(\mathcal{H}))$ =$=\{K(\cdot)+(\cdot)K^{*}+\Phi:K\in\mathcal{B}(\mathcal{H}),\Phi\in\mathsf{CP}(\mathcal{H})\}$
where $\mathfrak{u}(\mathcal{H})$ is the unitary algebra, i.e. the collection of all skew-adjoint bounded operators on $\mathcal{H}$.
Theorem 1.
Let an arbitrary separable complex Hilbert space $\mathcal{H}$ as well as $B\in\mathcal{B}^{1}(\mathcal{H})$ with ${\rm Re}({\rm tr}(B))\neq 0$ be given. Then
$\Xi_{B}:\{K\in\mathcal{B}(\mathcal{H}):{\rm Im}({\rm tr}(B^{*}K))=0\}\times\mathsf{CP}_{B}(\mathcal{H})$ $\to\mathsf{L}(\mathsf{CP}(\mathcal{H}))$
$(K,\Phi)$ $\mapsto K(\cdot)+(\cdot)K^{*}+\Phi$
is bijective.
Proof.
First we prove surjectivity. Starting from any $L\in\mathsf{L}(\mathsf{CP}(\mathcal{H}))$ one finds $K_{0}\in\mathcal{B}(\mathcal{H})$, $\Phi_{0}\in\mathsf{CP}(\mathcal{H})$ such that $L=K_{0}(\cdot)+(\cdot)K_{0}^{*}+\Phi_{0}$. Moreover, because $\Phi_{0}$ is completely positive and $\mathcal{H}$ is separable, as seen in Section 4 there exists $\{V_{j}\}_{j\in N}\subset\mathcal{B}(\mathcal{H})$, $N\subseteq\mathbb{N}$ such that $\Phi_{0}=\sum_{j\in N}V_{j}(\cdot)V_{j}^{*}$ and $\sum_{j\in N}V_{j}^{*}V_{j}$ is uniformly bounded. Then $v:=\{{\rm tr}(B^{*}V_{j})\}_{j\in N}\in\ell^{2}(N,\mathbb{C})$ (Lemma 6) so $\sum_{j\in N}(\frac{{\rm tr}(B^{*}V_{j})}{{\rm tr}(B^{*})})^{*}V_{j}$ converges strongly: for all $x\in\mathcal{H}\setminus\{0\}$:
$\sum_{j\in N}\big{\|}\tfrac{{\rm tr}(B^{*}V_{j})}{{\rm tr}(B^{*})}^{*}V_{j}x\big{\|}$ $=|{\rm tr}(B^{*})|^{-1}\sum_{j\in N}\big{|}{\rm tr}(B^{*}V_{j})\big{|}\,\|V_{j}x\|$
$\leq|{\rm tr}(B^{*})|^{-1}\Big{(}\sum_{j\in N}\big{|}{\rm tr}(B^{*}V_{j})\big{|}^{2}\Big{)}^{1/2}\Big{(}\sum_{j\in N}\big{\|}V_{j}x\big{\|}^{2}\Big{)}^{1/2}$
$=|{\rm tr}(B^{*})|^{-1}\|v\|_{2}\Big{(}\sum_{j\in N}\langle x,V_{j}^{*}V_{j}x\rangle\Big{)}^{1/2}$
$\leq|{\rm tr}(B^{*})|^{-1}\|v\|_{2}\Big{\|}\sum_{j\in N}V_{j}^{*}V_{j}\Big{\|}_{\infty}^{1/2}\|x\|$
Indeed, this computation together with Lemma 1.2.5 in Ringrose (1971) yields convergence of $\sum_{j\in N}\tfrac{{\rm tr}(B^{*}V_{j})}{{\rm tr}(B^{*})}^{*}V_{j}x$ as well as $\|\sum_{j\in N}(\frac{{\rm tr}(B^{*}V_{j})}{{\rm tr}(B^{*})})^{*}V_{j}\|_{\infty}\leq|{\rm tr}(B^{*})|^{-1}\|v\|_{2}\|\sum_{j\in N}V_{j}^{*}V_{j}\|_{\infty}^{1/2}$. Altogether this lets us define $\tilde{V}_{j}:=V_{j}-\frac{{\rm tr}(B^{*}V_{j})}{{\rm tr}(B^{*})}{\bf 1}\in\mathcal{B}(\mathcal{H})$ for all $j\in N$. Most importantly, $\sum_{j\in N}\tilde{V}_{j}^{*}\tilde{V}_{j}$ is uniformly bounded: for all $F\subseteq N$ finite
$\Big{\|}\sum_{j\in F}$ $\tilde{V}_{j}^{*}\tilde{V}_{j}\Big{\|}_{\infty}=\Big{\|}\sum_{j\in F}\Big{(}V_{j}-\frac{{\rm tr}(B^{*}V_{j})}{{\rm tr}(B^{*})}{\bf 1}\Big{)}^{*}\Big{(}V_{j}-\frac{{\rm tr}(B^{*}V_{j})}{{\rm tr}(B^{*})}{\bf 1}\Big{)}\Big{\|}_{\infty}$
$\leq\Big{\|}\sum_{j\in F}V_{j}^{*}V_{j}\Big{\|}_{\infty}+2\Big{\|}\sum_{j\in F}\Big{(}\frac{{\rm tr}(B^{*}V_{j})}{{\rm tr}(B^{*})}\Big{)}^{*}V_{j}\Big{\|}_{\infty}+\frac{\sum_{j\in F}|{\rm tr}(B^{*}V_{j})|^{2}}{|{\rm tr}(B)|^{2}}$
$\leq\Big{\|}\sum_{j\in N}V_{j}^{*}V_{j}\Big{\|}_{\infty}+2|{\rm tr}(B^{*})|^{-1}\|v\|_{2}\Big{\|}\sum_{j\in N}V_{j}^{*}V_{j}\Big{\|}_{\infty}^{1/2}+\frac{\sum_{j\in F}|{\rm tr}(B^{*}V_{j})|^{2}}{|{\rm tr}(B)|^{2}}$
This by Lemma 4 (ii) guarantees that $\Phi:=\sum_{j\in N}\tilde{V}_{j}(\cdot)\tilde{V}_{j}^{*}\in\mathsf{CP}(\mathcal{H})$ is well defined. Moreover, a straightforward computation shows $K_{0}(\cdot)+(\cdot)K_{0}^{*}+\Phi_{0}=\tilde{K}(\cdot)+(\cdot)\tilde{K}^{*}+\Phi$ when defining
$\tilde{K}:=K_{0}+\sum_{j\in N}\Big{(}\frac{{\rm tr}(B^{*}V_{j})}{{\rm tr}(B^{*})}\Big{)}^{*}V_{j}-\frac{\sum_{j\in N}|{\rm tr}(B^{*}V_{j})|^{2}}{2|{\rm tr}(B)|^{2}}{\bf 1}\,,$
cf. also (Davies, 1980, Eq. (1.4)). By definition, ${\rm tr}(B^{*}\tilde{V}_{j})=0$ for all $j$ so Lemma 6 shows $\Phi\in\mathsf{CP}_{B}(\mathcal{H})$. All that is left is to “shift” $\tilde{K}$ such that ${\rm Im}({\rm tr}(B^{*}\tilde{K}))=0$. Obviously, replacing $\tilde{K}$ by $\tilde{K}+i\lambda{\bf 1}$, $\lambda\in\mathbb{R}$ does not change $\tilde{K}(\cdot)+(\cdot)\tilde{K}^{*}$ and thus does not change $L$. For the shifted operator the trace condition then reads $0={\rm Im}({\rm tr}(B^{*}(\tilde{K}+i\lambda{\bf 1})))={\rm Im}({\rm tr}(B^{*}\tilde{K}))+\lambda{\rm Re}({\rm tr}(B^{*}))\,,$ so setting $K:=\tilde{K}-i\frac{{\rm Im}({\rm tr}(B^{*}\tilde{K}))}{{\rm Re}({\rm tr}(B))}{\bf 1}\in\mathcal{B}(\mathcal{H})$ yields $(K,\Phi)\in{\rm dom}(\Xi_{B})$ such that $\Xi_{B}(K,\Phi)=\tilde{K}(\cdot)+(\cdot)\tilde{K}^{*}+\Phi=K_{0}(\cdot)+(\cdot)K_{0}^{*}+\Phi_{0}=L$, as desired.
For injectivity, assume $K_{1}(\cdot)+(\cdot)K_{1}^{*}+\Phi_{1}=K_{2}(\cdot)+(\cdot)K_{2}^{*}+\Phi_{2}$ for some $K_{1},K_{2}\in\mathcal{B}(\mathcal{H})$ , $\Phi_{1},\Phi_{2}\in\mathsf{CP}_{B}(\mathcal{H})$ such that ${\rm Im}({\rm tr}(B^{*}K_{j}))=0$, $j=1,2$; equivalently, $(K_{2}-K_{1})(\cdot)+(\cdot)(K_{2}-K_{1})^{*}=\Phi_{1}-\Phi_{2}\in\mathsf{CP}_{B}(\mathcal{H})-\mathsf{CP}_{B}(\mathcal{H})$. Thus Proposition 2 shows that $K_{1}=K_{2}+i\lambda{\bf 1}$ for some $\lambda\in\mathbb{R}$. This has two consequences: On the one hand, this imaginary difference between $K_{1}$ and $K_{2}$ cancels in the sense that $K_{1}(\cdot)+(\cdot)K_{1}^{*}=K_{2}(\cdot)+(\cdot)K_{2}^{*}$ which in turn implies $\Phi_{1}=\Phi_{2}$. On the other hand, $K_{1}=K_{2}+i\lambda{\bf 1}$ together with the trace condition on $K_{1},K_{2}$ yields
$0={\rm Im}({\rm tr}(B^{*}K_{1}))={\rm Im}({\rm tr}(B^{*}K_{2}))+\lambda{\rm Re}({\rm tr}(B^{*}))=\lambda{\rm Re}({\rm tr}(B))\,.$
But ${\rm Re}({\rm tr}(B))\neq 0$ by assumption so $\lambda$ has to vanish, meaning $K_{1}=K_{2}$. This concludes the proof. ∎
Because $\mathsf{L}(\mathsf{CPTP}(\mathcal{H}))\subseteq\mathsf{L}(\mathsf{CP}(\mathcal{H}))$, as a direct consequence of Theorem 1 we obtain unique decompositions of generators of quantum-dynamical semigroups: this follows at once from the identification $K=-\frac{1}{2}\Phi^{*}({\bf 1})-iH$ (which is due to trace-preservation).
Corollary 1.
Let an arbitrary separable complex Hilbert space $\mathcal{H}$ as well as $B\in\mathcal{B}^{1}(\mathcal{H})$ with ${\rm Re}({\rm tr}(B))\neq 0$ be given. Then
$\hat{\Xi}_{B}:(H,\Phi)$ $\mapsto-i[H,\cdot]+\Phi-\Big{\{}\frac{\Phi^{*}({\bf 1})}{2},\cdot\Big{\}}\in\mathsf{L}(\mathsf{CPTP}(\mathcal{H}))$
with domain $\{(H,\Phi)\in i\mathfrak{u}(\mathcal{H})\times\mathsf{CP}_{B}(\mathcal{H}):\begin{subarray}{c}{\rm Im}({\rm tr}(\Phi(B)))=2\,{\rm Re}({\rm tr}(B^{*}H))\end{subarray}\}$ is bijective. In particular, if $B\in\mathcal{B}^{1}(\mathcal{H})$ is self-adjoint with ${\rm tr}(B)\neq 0$, then
$\hat{\Xi}_{B}:\{H\in i\mathfrak{u}(\mathcal{H}):{\rm tr}(BH)=0\}\times\mathsf{CP}_{B}(\mathcal{H})$ $\to\mathsf{L}(\mathsf{CPTP}(\mathcal{H}))$
$(H,\Phi)$ $\mapsto-i[H,\cdot]+\Phi-\Big{\{}\frac{\Phi^{*}({\bf 1})}{2},\cdot\Big{\}}$
is bijective.
From a physics perspective this corollary shows that any fixed reference state (more precisely: reference operator) gives rise to a unique splitting of GKSL-generators.
Finally a comment on how this result translates to the dual picture. There the object of interest is the collection of all $\Phi:\mathcal{B}(\mathcal{H})\to\mathcal{B}(\mathcal{Z})$ which are completely positive, unital (i.e. $\Phi({\bf 1}_{\mathcal{H}})={\bf 1}_{\mathcal{Z}}$), and ultraweakly continuous, denoted by $\mathsf{CPU}(\mathcal{H})_{\sigma}$. It is well known (and easy to verify) that $\Phi\in\mathsf{CPU}(\mathcal{H})_{\sigma}$ if and only if its dual map $\Phi^{*}$ is CPTP. Therefore $L\in\mathsf{L}(\mathsf{CPTP}(\mathcal{H}))$ is equivalent to $L^{*}\in\mathsf{L}(\mathsf{CPU}(\mathcal{H})_{\sigma})$ meaning the unique decomposition $(H,\Phi)$ of $L$ from Corollary 1 readily translates into a unique decomposition $(-H,\Phi^{*})$ of the dual generator $L^{*}$.
6 Open Questions
An obvious question right now is whether separability of the underlying Hilbert space is a necessary assumption in Theorem 1. The only point where this assumption was needed was Proposition 2 for which injectivity of the weighted Choi formalism (which only holds in the separable case, cf. Proposition 1) was crucial. This does of course not mean that our main result is wrong for non-separable Hilbert spaces, just that if it is true, then one needs an entirely different proof strategy for that.
Also one may wonder whether the reason the one-to-one relation between completely positive maps and positive semi-definite operators fails in infinite dimensions (Prop. 1) is that the Choi formalism was modified in an “unfavorable” way. Indeed, the “unweighted” Choi map $\Phi\mapsto\sum_{j,k\in J}|g_{j}\rangle\langle g_{k}|\otimes\Phi(|g_{j}\rangle\langle g_{k}|)$ establishes a correspondence between maps $:\mathcal{B}^{1}(\mathcal{H})\to\mathcal{B}^{1}(\mathcal{Z})$ of finite rank (i.e. maps of the form $X\mapsto\sum_{j=1}^{m}{\rm tr}(A_{j}X)B_{j}$ with $m<\infty$) and the algebraic tensor product $\mathcal{B}(\mathcal{H})\odot\mathcal{B}^{1}(\mathcal{Z})$. However, it is not clear what topology on the domain would allow for a completion of this correspondence to $\mathcal{B}(\mathcal{B}^{1}(\mathcal{H}),\mathcal{B}^{1}(\mathcal{Z}))$: the norm topology leads to a domain $\mathcal{K}(\mathcal{B}^{1}(\mathcal{H}),\mathcal{B}^{1}(\mathcal{Z}))$ which is too small, but something like the weak operator topology would likely be too weak and the domain would become too large.
Another question is concerned with the unbounded case, that is, the case of dynamical semigroups which are not norm- but only strongly continuous. While there are certain models where one can make sense of expressions like ${\rm tr}(B^{*}H)$ with $H$ unbounded (the relevant notion here are so-called “Schwartz operators”, cf. Keyl et al. (2016)), the bigger problem is that in the strongly continuous case there is no standard form of the corresponding generator anymore, refer to Siemon et al. (2017).
All questions posed above would, if solved, most likely require a vastly different set of tools as well as deep new insights into the Choi formalism or into generators of completely positive semigroups which is why we pointed them out explicitly.
Acknowledgements
I appreciate fruitful discussions with Federico Girotti at the “Workshop on dissipative evolutions on infinite dimensional and many-body quantum systems” 2023 in Nottingham. There it was pointed out to me that injectivity of $B$—which was an assumption in Prop. 2 ff. in an earlier draft—is not necessary for the main results to hold. Moreover, I am grateful to Sumeet Khatri for valuable comments on the Choi formalism in infinite dimensions. Moreover, we would like to thank the the anonymous referees for their helpful comments. This work has been supported by the Einstein Foundation (Einstein Research Unit on Quantum Devices) and the MATH+ Cluster of Excellence.
Appendix A Auxiliary Lemmata on Schatten Classes
The following result establishes how the strong operator topology interacts with the Schatten norms and is a generalization of (Widom, 1976, Prop. 2.1) from sequences to uniformly bounded nets, although the proof stays basically the same:
Lemma 9.
Let arbitrary complex Hilbert spaces $\mathcal{H},\mathcal{Z}$ and $A\in\mathcal{B}^{p}(\mathcal{H},\mathcal{Z})$ for $p\in[1,\infty]$, as well as nets $(B_{i})_{i\in I}\subset\mathcal{B}(\mathcal{H})$, $(C_{j})_{j\in J}\subset\mathcal{B}(\mathcal{Z})$ be given. If there exist bounded operators $B,C\in\mathcal{B}(\mathcal{H})$ such that $\|B_{i}x-Bx\|,\|C_{i}y-Cy\|\to 0$ for all $x,y\in\mathcal{H}$, and if there exists $\kappa>0$ such that $\|B_{i}\|_{\infty}\leq\kappa$, $\|C_{j}\|_{\infty}\leq\kappa$ for all $i\in I,j\in J$, then the net $(B_{i}AC_{j}^{*})_{i\in I,j\in J}\subset\mathcal{B}^{p}(\mathcal{H},\mathcal{Z})$ converges to $BAC^{*}$ in $p$-norm. If $I=J$, then $(B_{i}AC_{i}^{*})_{i\in I}\subset\mathcal{B}^{p}(\mathcal{H},\mathcal{Z})$ converges to $BAC^{*}$ in $p$-norm, as well. In particular this—even after dropping the above uniform boundedness requirement—holds for all sequences (i.e. $I,J\subseteq\mathbb{N}$) which strongly converge to some bounded operator.
Proof.
Consider any Schmidt decomposition $\sum_{k=1}^{\infty}s_{k}(A)|e_{k}\rangle\langle f_{k}|$ of the compact operator $A$ and let $\varepsilon>0$ be given. W.l.o.g. $\|B\|_{\infty},\|C\|_{\infty}<\kappa$—else define $\tilde{\kappa}:=\max\{\kappa,\|B\|_{\infty},\|C\|_{\infty}\}<\infty$. Now there exists $N\in\mathbb{N}$ such that
$\begin{cases}\sum_{k=N+1}^{\infty}s_{k}(A)^{p}<\frac{\varepsilon^{p}}{(3\kappa^{2})^{p}}&\text{ if }p\in[1,\infty)\\ s_{N+1}(A)<\frac{\varepsilon}{3\kappa^{2}}&\text{ if }p=\infty\end{cases}\,.$
Then, using the directed set property, strong convergence of $(B_{i})_{i\in I}$ implies the existence of $i_{0}\in I$ such that $\|B_{i}e_{k}-Be_{k}\|<\frac{\varepsilon}{6\kappa\sum_{k=1}^{N}s_{k}(A)}$ for all $i\succeq i_{0}$ and all $k=1,\ldots,N$; strong convergence of $(C_{j})_{j\in J}$ on the set $\{f_{1},\ldots,f_{N}\}$ yields a similar $j_{0}\in J$. Defining $A_{1}:=\sum_{k=1}^{N}s_{k}(A)|e_{k}\rangle\langle f_{k}|$ and $A_{2}:=A-A_{1}$ we for all $j\succeq(i_{0},j_{0})$ (i.e. $i\succeq i_{0}$ and $j\succeq j_{0}$) compute
[…]
$\leq\|B_{i}A_{1}C_{j}^{*}-B_{i}A_{1}C_{j}^{*}\|_{p}+\|B_{i}A_{2}C_{j}^{*}\|_{p}+\|BA_{2}C^{*}\|_{p}$
$\leq\Big{\|}\sum_{k=1}^{N}s_{k}(A)\big{(}|B_{i}e_{k}\rangle\langle C_{j}f_{k}|-|Be_{k}\rangle\langle Cf_{k}|\big{)}\Big{\|}_{p}+2\kappa^{2}\|A_{2}\|_{p}$
$<\Big{\|}\sum_{k=1}^{N}s_{k}(A)|(B_{i}-B)e_{k}\rangle\langle C_{j}f_{k}|\Big{\|}_{p}+\Big{\|}\sum_{k=1}^{N}s_{k}(A)|Be_{k}\rangle\langle(C_{j}-C)f_{k}|\Big{\|}_{p}+\frac{2\varepsilon}{3}$
$\leq\sum_{k=1}^{N}s_{k}(A)\|(B_{i}-B)e_{k}\|\|C_{j}f_{k}\|+\sum_{k=1}^{N}s_{k}(A)\|Be_{k}\|\|(C_{j}-C)f_{k}\|+\frac{2\varepsilon}{3}$
$\leq\kappa\sum_{k=1}^{N}s_{k}(A)\big{(}\|B_{i}e_{k}-Be_{k}\|+\|C_{j}f_{k}-Cf_{k}\|\big{)}+\frac{2\varepsilon}{3}<\varepsilon\,.$
If $I=J$ then one shows $\|B_{i}AC_{i}^{*}-BAC^{*}\|_{p}\to 0$ analogously. Now the additional statement about sequences of operators follows at once from the uniform boundedness principle. ∎
This has an immediate consequence for block approximations of Schatten class as well as for general bounded operators:
Corollary 2.
Let arbitrary complex Hilbert spaces $\mathcal{H},\mathcal{Z}$ and $A\in\mathcal{L}(\mathcal{H},\mathcal{Z})$ be given. For any orthonormal bases $\{f_{k}\}_{k\in K}$, $\{g_{j}\}_{j\in J}$ of $\mathcal{Z}$, $\mathcal{H}$, respectively, as well as any finite subsets $J^{\prime}\subseteq J$, $\mathcal{K}^{\prime}\subseteq K$ define
$A_{J^{\prime},K^{\prime}}:=\sum_{j\in J^{\prime}}\sum_{k\in K^{\prime}}\langle f_{k},Ag_{j}\rangle|f_{k}\rangle\langle g_{j}|\in\mathcal{B}(\mathcal{H},\mathcal{Z})\,.$
Then the following statements hold.
  • (i)
    If $A\in\mathcal{B}^{p}(\mathcal{H})$ for some $p\in[1,\infty]$, then $(A_{J^{\prime},K^{\prime}})_{J^{\prime}\subseteq J,\mathcal{K}^{\prime}\subseteq K\text{ finite}}$ converges to $A$ in $p$-norm.
  • (ii)
    If $A$ is bounded, then $(A_{J^{\prime},K^{\prime}})_{J^{\prime}\subseteq J,\mathcal{K}^{\prime}\subseteq K\text{ finite}}$ converges to $A$ in the strong operator topology.
In particular, for all orthonormal bases $\{f_{k}\}_{k\in K}$, $\{g_{j}\}_{j\in J}$ of $\mathcal{Z}$, $\mathcal{H}$, respectively, ${\rm span}\{|f_{k}\rangle\langle g_{j}|:j\in J,k\in K\}$ is dense in $(\mathcal{B}^{p}(\mathcal{H},\mathcal{Z}),\|\cdot\|_{p})$ for all $p\in[1,\infty]$.
Proof.
(i): Given any finite subsets $J^{\prime}\subseteq J$, $K^{\prime}\subseteq K$, respectively, define the orthogonal projections $\Pi_{J^{\prime}}:=\sum_{j\in J^{\prime}}|g_{j}\rangle\langle g_{j}|$, $\tilde{\Pi}_{K^{\prime}}:=\sum_{k\in K^{\prime}}|f_{k}\rangle\langle f_{k}|$. The corresponding nets $(\Pi_{J^{\prime}})_{J^{\prime}\subseteq J\text{ finite}}$, $(\tilde{\Pi}_{K^{\prime}})_{K^{\prime}\subseteq K\text{ finite}}$ are well known to converge to ${\bf 1}_{\mathcal{H}}$, ${\bf 1}_{\mathcal{Z}}$, respectively, in the strong operator topology. With this, Lemma 9 shows $\|A_{J^{\prime},K^{\prime}}-A\|_{p}=\|\tilde{\Pi}_{K^{\prime}}A\Pi_{J^{\prime}}-A\|_{p}\to 0$ because $\|\Pi_{J^{\prime}}\|_{\infty}=\|\tilde{\Pi}_{K^{\prime}}\|_{\infty}=1$ for all $J^{\prime}\subseteq J,K^{\prime}\subseteq K$. The additional statement also follows from this. (ii): As before one for all $x\in\mathcal{H}$ computes
$\|(A_{J^{\prime},K^{\prime}}-A)x\|$ $=\|\tilde{\Pi}_{K^{\prime}}A\Pi_{J^{\prime}}x-Ax\|$
$\leq\|A\|_{\infty}\|(\Pi_{J^{\prime}}-{\bf 1}_{\mathcal{H}})x\|+\|(\tilde{\Pi}_{K^{\prime}}-{\bf 1}_{\mathcal{Z}})Ax\|\to 0\,.\qed$
Finally, we prove a technical lemma about trace norm convergence of certain sums of operators.
Lemma 10.
Let complex Hilbert spaces $\mathcal{H},\mathcal{Z}$ and $\{V_{j}\}_{j\in J}\subset\mathcal{B}(\mathcal{H},\mathcal{Z})$ be given such that $\{\sum_{j\in F}V_{j}^{*}V_{j}\}_{F\subseteq J\text{ finite}}$ is uniformly bounded. Then for all $A\in\mathcal{B}^{1}(\mathcal{H})$ positive semi-definite $\{\sum_{j\in F}V_{j}AV_{j}^{*}\}_{F\subseteq J\text{ finite}}$ converges in trace norm. Moreover, if $X\in\mathcal{B}(\mathcal{H})$ denotes the limit of $\{\sum_{j\in F}V_{j}^{*}V_{j}\}_{F\subseteq J\text{ finite}}$, then ${\rm tr}(\sum_{j\in J}V_{j}AV_{j}^{*})={\rm tr}(XA)$.
Proof.
Given $A\in\mathcal{B}^{1}(\mathcal{H})$ positive semi-definite (i.e. $A=\sum_{k\in N}s_{k}(A)|g_{k}\rangle\langle g_{k}|$ for some orthonormal system $\{g_{k}\}_{k\in N}$ in $\mathcal{H}$) our goal is to show that for all $\varepsilon>0$ there exists $F_{\varepsilon}\subseteq J$ finite such that $\|\sum_{j\in F}V_{j}AV_{j}^{*}\|_{1}<\varepsilon$ for all $F\subseteq J\setminus F_{\varepsilon}$ finite (this is sufficient due to Lemma 1.2.2 in Ringrose (1971)). By Lemma 3, $\{\sum_{j\in F}V_{j}^{*}V_{j}\}_{F\subseteq J\text{ finite}}$ converges strongly and ultraweakly—so in particular weakly—to some $X\in\mathcal{B}(\mathcal{H})$. W.l.o.g. $X\neq 0$ (else $V_{j}=0$ for all $j\in J$) as well as $A\neq 0$.
Now let any $\varepsilon>0$ be given. Because $(s_{k}(A))_{k\in N}\in\ell^{1}(N,\mathbb{C})$, there exists $N_{\varepsilon}\subseteq N$ such that $N\setminus N_{\varepsilon}$ is finite and that $\sum_{k\in N\setminus N_{\varepsilon}}s_{k}(A)>0$, as well as $\sum_{k\in N_{\varepsilon}}s_{k}(A)<\frac{\varepsilon}{2\|X\|_{\infty}}$. Moreover, because $\{\sum_{j\in F}V_{j}^{*}V_{j}\}_{F\subseteq J\text{ finite}}$ converges weakly, $\{\langle g_{k},\sum_{j\in F}V_{j}^{*}V_{j}g_{k}\rangle\}_{F\subseteq J\text{ finite}}=\{\sum_{j\in F}\|V_{j}g_{k}\|^{2}\}_{F\subseteq J\text{ finite}}$ is summable for all $k\in N\setminus N_{\varepsilon}$. This way one iteratively finds $F_{\varepsilon}\subseteq J$ finite such that $\sum_{j\in F}\|V_{j}g_{k}\|^{2}<\frac{\varepsilon}{2\sum_{k\in N\setminus N_{\varepsilon}}s_{k}(A)}$ for all $F\subseteq J\setminus F_{\varepsilon}$ finite and all $k\in N\setminus N_{\varepsilon}$. This is all we need in order to verify the above Cauchy criterion: for all $F\subseteq J\setminus F_{\varepsilon}$ finite, using the triangle inequality we can upper bound $\|\sum_{j\in F}V_{j}AV_{j}^{*}\|_{1}$ via
$\sum_{j\in F}$ $\sum_{k\in N\setminus N_{\varepsilon}}s_{k}(A)\big{\|}V_{j}|g_{k}\rangle\langle g_{k}|V_{j}^{*}\big{\|}_{1}+\sum_{j\in F}\sum_{k\in N_{\varepsilon}}s_{k}(A)\big{\|}V_{j}|g_{k}\rangle\langle g_{k}|V_{j}^{*}\big{\|}_{1}$
$\leq\sum_{k\in N\setminus N_{\varepsilon}}s_{k}(A)\Big{(}\sum_{j\in F}\|V_{j}g_{k}\|^{2}\Big{)}+\sum_{k\in N_{\varepsilon}}s_{k}(A)\Big{\langle}g_{k},\sum_{j\in F}V_{j}^{*}V_{j}g_{k}\Big{\rangle}$
$<\frac{\varepsilon}{2}+\sum_{k\in N_{\varepsilon}}s_{k}(A)\Big{\|}\sum_{j\in F}V_{j}^{*}V_{j}\Big{\|}_{\infty}\leq\frac{\varepsilon}{2}+\sum_{k\in N_{\varepsilon}}s_{k}(A)\|X\|_{\infty}<\frac{\varepsilon}{2}+\frac{\varepsilon}{2}=\varepsilon\,.$
The final claim follows from the trace-norm convergence we just showed, together with ultraweak convergence of $\{\sum_{j\in F}V_{j}^{*}V_{j}\}_{F\subseteq J\text{ finite}}$:
${\rm tr}\Big{(}\sum_{j\in J}V_{j}AV_{j}^{*}\Big{)}=\lim_{F}{\rm tr}\Big{(}\sum_{j\in F}V_{j}AV_{j}^{*}\Big{)}=\lim_{F}{\rm tr}\Big{(}\sum_{j\in F}V_{j}^{*}V_{j}A\Big{)}={\rm tr}(XA)$